### v2

# Weighted heat kernel comparison theorems and its applications in spectral geometry

Abstract
In this paper, we firstly establish weighted heat kernel comparison theorems for the weighted heat equation on complete manifolds with radial curvatures bounded, and then by mainly using this conclusion, we can obtain two eigenvalue comparison theorems for the first Dirichlet eigenvalue of the Witten-Laplacian as applications in spectral geometry.
Faculty of Mathematics and Statistics,
Key Laboratory of Applied Mathematics of Hubei Province,
Hubei University, Wuhan 430062, China
Key Laboratory of Intelligent Sensing System and Security (Hubei University), Ministry of Education
Email: jiner120@163.com
1 Introduction
Given an $n$-dimensional Riemannian manifold $M^{n}$ with the metric $\langle\cdot,\cdot\rangle$, assume that $\phi\in C^{\infty}(M^{n})$ is a smooth111 One might find that it is sufficient to get some conclusions in this paper, provided $\phi$ is only $C^{2}$. However, for the convenience and in order to avoid any potential confusion, we assume that $\phi$ is smooth and wish to ignore the discussion on seeking the weakest regularity of the potential function $\phi$ such that all the assertions in this paper are still valid, which of course is not our main purpose here. real-valued function defined on $M^{n}$, and $\Omega\subset M^{n}$ is a bounded domain. Then it is natural to define the notion of smooth metric measure space $(M^{n},\langle\cdot,\cdot\rangle,e^{-\phi}dv)$, which actually endows the Riemannian volume element $dv$ of $M^{n}$ with a conformal factor $e^{-\phi}$. On the smooth metric measure space $(M^{n},\langle\cdot,\cdot\rangle,e^{-\phi}dv)$, the Witten-Laplacian (also called weighted Laplacian, drifting Laplacian, or $\phi$-Laplacian) and the $N$-dimensional Bakry-Émery Ricci curvature tensor can be well-defined as follows
$\Delta_{\phi}:=\Delta-\langle\nabla\cdot,\nabla\phi\rangle=e^{\phi}\mathrm{div}(e^{-\phi}\nabla\cdot)$
and
$\mathrm{Ric}^{N}_{\phi}:=\mathrm{Ric}+\mathrm{Hess}\phi-\frac{d\phi\otimes d\phi}{N-n},$
where $\Delta$, $\nabla$ denote the Laplace and the gradient operators on $M^{n}$ respectively, $\mathrm{div}$ is the divergence operator on $M^{n}$, $\mathrm{Ric}$ is the Ricci tensor on $M^{n}$, and $\mathrm{Hess}$ is the Hessian operator on $M^{n}$ associated to the metric $\langle\cdot,\cdot\rangle$. For the $N$-dimensional Bakry-Émery Ricci curvature $\mathrm{Ric}^{N}_{\phi}$, one knows $N>n$ or $N=n$ if $\phi$ is a constant function. Specially, when $N=\infty$, one can define the so-called $\infty$-dimensional Bakry-Émery Ricci curvature (simply, Bakry-Émery Ricci curvature or weighted Ricci curvature) as follows
$\mathrm{Ric}_{\phi}=\mathrm{Ric}+\mathrm{Hess}\phi.$
D. Bakry and M. Émery [4] introduced these curvature tensors, and studied their relationship with diffusion processes. For the domain $\Omega$ on $(M^{n},\langle\cdot,\cdot\rangle,e^{-\phi}dv)$, one can also define a notion weighted volume (or $\phi$-volume) as follows222 Generally, except specified, same mathematical notations would have the same meanings in this paper.
$\mathrm{vol}_{\phi}(\Omega)=|\Omega|_{n,\phi}:=\int_{\Omega}d\mu=\int_{\Omega}e^{-\phi}dv,$
where $d\mu:=e^{-\phi}dv$ stands for the weighted volume density. Clearly, if $\phi=const.$ is a constant function, the Witten-Laplacian $\Delta_{\phi}$, Bakry-Émery Ricci curvature $\mathrm{Ric}_{\phi}$, the weighted volume $|\Omega|_{n,\phi}$ would degenerate into the usual Laplacian $\Delta$, the Ricci curvature tensor $\mathrm{Ric}$, and the volume $\mathrm{vol}(\Omega)=|\Omega|_{n}=\int_{\Omega}dv$ (i.e. the volume of $\Omega$), respectively. We wish to refer readers to [5, 46, 47] for further details about geometric properties of the Bakry-Émery Ricci curvature tensor.
In the past four decades, smooth metric measure spaces have been extensively investigated in geometric analysis, and have close relation with probability theory and optimal transport. We wish to mention again several important examples listed in our previous work [72, Section 1] to support this viewpoint — G. Perelman’s pioneering work [60] on the entropy formula of the Ricci flow under a weighted measure, Munteanu-Wang’s series works [57, 58, 59] about geometry of manifolds with weighted densities, Wei-Wylie’s interesting work [68] on some comparison theorems for manifolds with Bakry-Émery Ricci curvature bounded. Besides, recently, by mainly using the rearrangement technique and suitably constructing trial functions, under the constraint of fixed weighted volume, R. F. Chen and J. Mao [21] successfully obtain several isoperimetric inequalities for the first and the second Dirichlet eigenvalues, the first nonzero Neumann eigenvalue of the Witten Laplacian on bounded domains in space forms. These spectral isoperimetric inequalities extend the classical ones (i.e., the Faber-Krahn inequality, the Hong-Krahn-Szegő inequality, and the Szegő-Weinberger inequality) of the Laplacian.
A smooth metric measure metric space $(M^{n},\langle\cdot,\cdot\rangle,e^{-\phi}dv)$ is said to be quasi-Einstein if
$\mathrm{Ric}^{N}_{\phi}=\tau\langle\cdot,\cdot\rangle$
for some constant $\tau$. Specially, when $N=\infty$, the above equality becomes
$\mathrm{Ric}_{\phi}=\tau\langle\cdot,\cdot\rangle,$
and in this setting, $(M^{n},\langle\cdot,\cdot\rangle,e^{-\phi}dv)$ is called a gradient Ricci soliton. A gradient Ricci soliton is called expanding, steady or shrinking if $\tau<0$, $\tau=0$, and $\tau>0$, respectively. Ricci solitons are natural extension of Einstein manifolds, and play an important role in the singularity analysis of Ricci flow (see e.g. [6, 60]). We wish to refer readers to a nice survey [13] on the topic of Ricci solitons.
On the smooth metric measure metric space $(M^{n},\langle\cdot,\cdot\rangle,e^{-\phi}dv)$, it is natural to consider the following weighted heat equation333 For convenience, usually the partial derivative $\partial u/\partial t$ of $u(x,t)$ was rewritten as $u_{t}$. Similarly, $\partial u/\partial x$, $\partial^{2}u/\partial x^{2}$ can be rewritten as $u_{x}$ and $u_{xx}$, respectively. This convention on rewriting partial derivatives of multivariable functions would also be used in the sequel.
$\Delta_{\phi}u-\frac{\partial u}{\partial t}=0$ (1.1)
instead of the classical heat equation, and correspondingly, the weighted heat operator $\mathcal{L}^{\phi}$ can be well-defined as follows
$\mathcal{L}^{\phi}:=\Delta_{\phi}-\frac{\partial}{\partial t}.$
Obviously, if $\phi=const.$, then the weighted heat equation (1.1), the weighted heat operator $\mathcal{L}^{\phi}$ degenerate into the heat equation $\Delta u-u_{t}=0$ and the heat operator $\mathcal{L}:=\Delta-\partial/\partial t$, respectively. Besides, if a solution $u$ to the equation (1.1) is independent of the time variable $t$, then $u$ is a weighted harmonic function (also called $\phi$-harmonic function).
We need the following notion:
Definition 1.1.
Let $M^{n}$ be a Riemannian $n$-manifold and $\phi$ be a real-valued smooth function on $M^{n}$. A fundamental solution444 It is easy to see that if $u$ is a solution to the equation (1.1), then u-u is also. Hence, in some literatures the positivity of the heat kernel has not been specially mentioned. However, as a default, the kernel is usually required to be positive. We shall not mention this point again in the sequel., which is called the weighted heat kernel, of the weighted heat equation (1.1) on $M^{n}$ is a continuous function $H^{\phi}=H^{\phi}(x,y,t)$, defined on $M^{n}\times M^{n}\times(0,\infty)$, which is $C^{2}$ with respect to $x$, $C^{1}$ with respect to $t$, and which satisfies
$\mathcal{L}^{\phi}_{x}H^{\phi}=0,\qquad\lim\limits_{t\rightarrow 0}H^{\phi}(x,y,t)=\delta^{\phi}_{y}(x),$
where $\delta^{\phi}_{y}(x)$ is the weighted Dirac delta function, that is, for all smooth functions $f$ on $M^{n}$ with compact support (i.e. $f\in C^{\infty}_{0}(M^{n})$), we have, for every $y\in M^{n}$,
$\lim\limits_{t\rightarrow 0}\int\limits_{M^{n}}f(x)H^{\phi}(x,y,t)e^{-\phi}dv(x)=f(y).$
Remark 1.2.
In some literatures, see e.g. [70, 71], the weighted heat equation (1.1), the weighted heat kernel are also called $\phi$-heat equation and $\phi$-heat kernel, respectively. Clearly, if $\phi=const.$, then the weighted heat kernel reduces into the usual heat kernel of the heat equation.
It is well-known that the heat kernel of the heat equation in the Euclidean $n$-space $\mathbb{R}^{n}$ is explicitly given as follows
$H(x,y,t)=\frac{1}{(4\pi t)^{n/2}}\times\exp\left(-\frac{|x-y|^{2}}{4t}\right).$
Now, we wish to recall several examples to intuitively show that even in the one dimensional setting, the weighted heat kernel in several nontrivial cases (i.e. $\phi$ is not a constant function) is more complicated than the one in the trivial case, i.e. $\phi=const.$ and $H^{\phi}(x,y,t)=H(x,y,t)=\frac{1}{(4\pi t)^{1/2}}\times\exp\left(-\frac{|x-y|^{2}}{4t}\right)$ corresponding to the trivial case.
Example 1.3 ($\phi$-heat kernel for steady Gaussian soliton [70]).
Denote by $(\mathbb{R},g_{E},e^{-\phi}dx)$ a $1$-dimensional steady Gaussian soliton, where $g_{E}$ is the Euclidean metric and $\phi(x)=\pm x$. Then $\mathrm{Ric}_{\phi}=0$, $\Delta_{\phi}=\frac{d^{2}}{dx^{2}}\mp\frac{d}{dx}$, and by using the separation of variables method, the weighted heat kernel is given by
$H^{\phi}(x,y,t)=\frac{e^{\pm\frac{x+y}{2}}\cdot e^{-t/4}}{(4\pi t)^{1/2}}\times\exp\left(-\frac{|x-y|^{2}}{4t}\right).$
For detailed calculations, see [70, Appendix].
Example 1.4 (Mehler heat kernel for shrinking Gaussian soliton [38]).
Let $(\mathbb{R},g_{E},e^{-\phi}dx)$ be a $1$-dimensional shrinking Gaussian soliton, where $g_{E}$ is the Euclidean metric and $\phi(x)=x^{2}$. Then $\mathrm{Ric}_{\phi}=2$, $\Delta_{\phi}=\frac{d^{2}}{dx^{2}}-2x\frac{d}{dx}$, and in this setting the weighted heat kernel is given by
$H^{\phi}(x,y,t)=\frac{1}{(2\pi\sinh 2t)^{1/2}}\times\exp\left(\frac{2xye^{-2t}-(x^{2}+y^{2})e^{-4t}}{1-e^{-4t}}+t\right).$
Example 1.5 (Mehler heat kernel for expanding Gaussian soliton [38]).
Let $(\mathbb{R},g_{E},e^{-\phi}dx)$ be a $1$-dimensional expanding Gaussian soliton, where $g_{E}$ is the Euclidean metric and $\phi(x)=-x^{2}$. Then $\mathrm{Ric}_{\phi}=-2$, $\Delta_{\phi}=\frac{d^{2}}{dx^{2}}+2x\frac{d}{dx}$, and in this setting the weighted heat kernel is given by
$H^{\phi}(x,y,t)=\frac{1}{(2\pi\sinh 2t)^{1/2}}\times\exp\left(\frac{2xye^{-2t}-(x^{2}+y^{2})}{1-e^{-4t}}-t\right).$
There are so many mathematicians having interest on the (non-weighted or weighted) heat kernel and its applications, and many interesting results have been obtained. Here we wish to give a partial list of those results which influence me a lot. In 1981, S. Y. Cheng, P. Li and S. T. Yau [26] gave uniform Gaussian estimates for the heat kernel on Riemannian manifolds with sectional curvature bounded from below. Their estimates were improved by J. Cheeger, M. Gromov, M. Taylor [17] to manifolds with bounded geometry. By using the gradient estimate technique and the Harnack inequality, P. Li and S. T. Yau [44] obtained sharp Gaussian upper and lower bounds for the Schrödinger operator on Riemannian manifolds with nonnegative Ricci curvature. For smooth metric measure spaces with $\mathrm{Ric}^{N}_{\phi}$ ($N<\infty$) bounded from below by a negative quadratic function, X. D. Li [43] got Gaussian estimates for the weighted heat kernel, and proved an $L^{1}_{\phi}$-Liouville theorem. In [69], J. Y. Wu obtained weighted heat kernel estimates on smooth metric measure spaces with the weighted Ricci curvature bounded from below by a negative constant and the potential function bounded. By using the De Giorgi-Nash-Moser theory and the weighted version of Davies’s integral estimate [27], J. Y. Wu and P. Wu [70, Theorem 1.1] successfully gave a local Gaussian upper bound for the weighted heat kernel on complete noncompact smooth metric measure spaces $(M^{n},\langle\cdot,\cdot\rangle,e^{-\phi}dv)$ with $\mathrm{Ric}_{\phi}\geq 0$. Their upper bound estimate generalizes the classical result of Li-Yau [44]. As applications of their local Gaussian upper bound estimate for the weighted heat kernel, they showed that if furthermore the potential function $\phi$ has at most quadratic growth, then a sharp $L^{1}_{\phi}$-Liouville theorem for $\phi$-subharmonic functions and an $L^{1}_{\phi}$-uniqueness property for nonnegative solutions of the weighted heat equation can be obtained (see [70, Theorem 1.5]). Besides, it is not hard to see from [14, 40] that any complete noncompact shrinking or steady gradient Ricci soliton has nonnegative weighted Ricci curvature and its potential function $\phi$ should have at most quadratic growth. Then together with Wu-Wu’s $L^{1}_{\phi}$-Liouville theorem mentioned above, it is easy to know that any $L^{1}_{\phi}$-integral $\phi$-subharmonic function on gradient shrinking and steady Ricci solitons must be constant (see [70, Corollary 1.6]). Later, as a sequel of the work [70], by mainly using the De Giorgi-Nash-Moser theory, J. Y. Wu and P. Wu [71, Theorem 1.1] obtained local Gaussian (upper and lower) bounds for the weighted heat kernel on complete noncompact smooth metric measure spaces $(M^{n},\langle\cdot,\cdot\rangle,e^{-\phi}dv)$ with Bakry-Émery Ricci curvature bounded from below by some nonnegative constant (i.e. $\mathrm{Ric}_{\phi}\geq-(n-1)K$ for some constant $K\geq 0$). Moreover, when the curvature assumption was reduced to $\mathrm{Ric}_{\phi}\geq 0$ (i.e. $K=0$), they pointed out that the local Gaussian lower bound for the weighted heat kernel can be achieved by the $\phi$-heat kernel for steady Gaussian soliton (explicitly given in Example 1.3) as long as the time $t$ is large enough. Besides, as applications, they used their local Gaussian (upper and lower) bounds for the weighted heat kernel to obtain estimates for closed eigenvalues and for $\phi$-Green’s function of the Witten-Laplacian – see [71, Sections 6 and 7] for details. In our recent work [72], by utilizing some functional inequalities and combining with De Giorgi-Moser iteration technique, some estimates for the weighted $L^{\infty}$-norm of eigenfunctions of the Witten-Laplacian on smooth metric measure spaces have been given. Besides, we also proved the regularity and uniqueness of the weighted heat kernel on smooth metric measure spaces, which generalized Chavel-Feldman’s related results in [15].
In this paper, we wish to give a comparison theorem for the weighted heat kernel (of the weighted heat equation) on complete manifolds with radial curvatures bounded. In this setting, the model space would be spherically symmetric manifolds. In order to state our main result clearly, we need to recall some preliminary knowledge about the exponential mapping, the cut locus, the radial direction, the injectivity radius, the radial (Ricci or sectional) curvature having a bound given by a continuous function (see Definitions 2.2 and 2.3), spherically symmetric manifolds (see Definition 2.1), and so on. However, in order to control the length of this section, we wish to give preliminaries in Section 2 of this paper. By the way, these preliminaries can also be found in our previous works [33, Section 1], [48, Section 2.1 of Chapter 2], [49, Section 2].
For a given Riemannian $n$-manifold $M^{n}$ with radial (Ricci or sectional) curvature bounded by a continuous function $\kappa(s)$ of the distance parameter $s:=d_{M^{n}}(q,\cdot)$, $q\in M^{n}$, where $d_{M^{n}}$ denotes the distance function on $M^{n}$, we consider the following initial value problem
$\left\{\begin{array}[]{lll}f^{\prime\prime}(s)+\kappa(s)f(s)=0,\qquad&0<s<l,\\ f^{\prime}(0)=1,~f(0)=0,\\ f(s)>0,\qquad&0<s<l,\end{array}\right$. (1.5)
which will be used to determine our model space $0,l)\times_{f}\mathbb{S}^{n-1}$ (i.e. spherically symmetric manifolds, which are called generalized space forms in [42]). See Definition 2.1 in Section 2 for the precise statement such that a domain to be spherically symmetric. Readers can check our previous work [33, pp. 705–709] for a brief introduction and an interesting spectral asymptotic property of spherically symmetric manifolds. Obviously, if in (1.5), $\kappa(s)\equiv K$ for some constant $K$, then
$f(s)=\left\{\begin{array}[]{llll}\frac{\sin\sqrt{K}s}{\sqrt{K}},&\quad l=\frac{\pi}{\sqrt{K}}&\quad K>0,\\ s,&\quad l=+\infty&\quad K=0,\\ \frac{\sinh\sqrt{-K}s}{\sqrt{-K}},&\quad l=+\infty&\quad K<0,\end{array}\right$.
and correspondingly, the model space reduces to space forms.
We need to require:
Property 1. The potential function $\phi:(0,l)\rightarrow\mathbb{R}$ is additionally required to be a radial function with respect to the Riemannian distance parameter, and moreover, its smoothness at the starting point (no matter on the given complete manifold but also on the model space) should be assured.
Our main result is:
Theorem 1.6.
Given a complete Riemannian $n$-manifold $M^{n}$, $n\geq 2$, and assuming that the potential function $\phi$ satisfies Property 1, we can obtain:
(1) if $M^{n}$ has a radial Ricci curvature lower bound $(n-1)\kappa(s)$ w.r.t. some point $q\in M^{n}$, then, for $r_{0}<\min\{\ell(q),l\}$, the inequality
$H^{\phi}(q,y,t)\geq H^{\phi}_{-}\left(d_{M^{-}}\left(q^{-},z\right),t\right),$ (1.7)
holds for all $(y,t)\in B(q,r_{0})\times(0,\infty)$ with $d_{M^{n}}(q,y)=d_{M^{-}}(q^{-},z)$ for any $z\in M^{-}$, where $\ell(q)$ is defined by555 In fact, if one checks the details in Section 2, he or she would find that the quantity $\ell(q)$ is closely related with the first conjugate points of $q$ along different unit minimizing geodesics $\gamma_{\xi}(s)=\exp_{q}(s\xi)$, with $\xi\in S_{q}M^{n}\subset T_{q}M^{n}$. (2.7), the model space is $M^{-}=:[0,l)\times_{f}\mathbb{S}^{n-1}$ with the base point $q^{-}$ and $f$ determined by (1.5), $d_{M^{-}}$ denotes the distance function on $M^{-}$. Moreover, the equality in (1.7) holds at some $(y_{0},t_{0})\in B(q,r_{0})\times(0,\infty)$ if and only if $B(q,r_{0})$ is isometric to $\mathscr{B}_{n}(q^{-},r_{0})$. Here $B(q,r_{0})$ is the geodesic ball on $M^{n}$ with center $q$ and radius $r_{0}$, while $\mathscr{B}_{n}(q^{-},r_{0})$ is the geodesic ball on $M^{-}$ with center $q^{-}$ and the same radius $r_{0}$;
(2) if $M^{n}$ has a radial sectional curvature upper bound $\kappa(s)$ w.r.t. $q\in M^{n}$, then, for $r_{0}<\min\{\mathrm{inj}(q),l\}$, the inequality
$H^{\phi}(q,y,t)\leq H^{\phi}_{+}\left(d_{M^{+}}\left(q^{+},z\right),t\right),$ (1.8)
holds for all $(y,t)\in B(q,r_{0})\times(0,\infty)$ with $d_{M^{n}}(q,y)=d_{M^{+}}(q^{+},z)$ for any $z\in M^{+}$, where $\mathrm{inj}(q)$ defined by (2.5) is the injectivity radius at $q$, the model space is $M^{+}=:[0,l)\times_{f}\mathbb{S}^{n-1}$ with the base point $q^{+}$ and $f$ determined by (1.5), $d_{M^{+}}$ denotes the distance function on $M^{+}$. Moreover, the equality in (1.8) holds at some $(y_{0},t_{0})\in B(q,r_{0})\times(0,\infty)$ if and only if $B(q,r_{0})$ is isometric to $\mathscr{B}_{n}(q^{+},r_{0})$. Here $\mathscr{B}_{n}(q^{+},r_{0})$ denotes the geodesic ball on $M^{+}$ with center $q^{+}$ and radius $r_{0}$.
(The boundary condition will either be Dirichlet or Neumann.)
Example 1.7 ($1$-dimensional model).
Let $(\mathbb{R},g_{E})$ be the Euclidean $1$-space with the Euclidean metric $g_{E}$. For a fixed point $x_{0}\in\mathbb{R}$, set $r:=g_{E}(x_{0},y)=|y-x_{0}|$, which of course measures the Euclidean distance between two points $x_{0}$, $y$ in $\mathbb{R}$. If the potential function $\phi$ was chosen to be $\phi=\phi(r)=\pm r$, which is radial with respect to $x_{0}$, then $\mathrm{Ric}_{\phi}=0$, $\Delta_{\phi}=\frac{d^{2}}{dr^{2}}\mp\frac{d}{dr}$, and by using the separation of variables method (see Appendix A), the weighted heat kernel is given by
$H^{\phi}(x_{0},y,t)=H^{\phi}(r,t)$ $=$ $\frac{e^{\pm r}\cdot e^{-t/4}}{(4\pi t)^{1/2}}\times\exp\left(-\frac{r^{2}}{4t}\right)$
$=$ $\frac{e^{\pm|y-x_{0}|}\cdot e^{-t/4}}{(4\pi t)^{1/2}}\times\exp\left(-\frac{|y-x_{0}|^{2}}{4t}\right).$
Remark 1.8.
In fact, from Example 1.7 above, it is not hard to explicitly get the weighted heat kernel (with a prescribed Dirichlet or Neumann boundary condition) on balls in $\mathbb{R}^{n}$ if the potential function $\phi$ was chosen to be $\phi(r)=\pm r$. Of course, in this situation, $\phi$ is radial with respect to the center of the given Euclidean $n$-ball.
Example 1.9.
(see [38, Subsection 4.6]) Let $\mathbb{H}^{3}$ be the hyperbolic $3$-space with constant666 For the weighted heat kernel $H^{\phi}$ shown in Example 1.9, there is no essential difference between $\mathbb{H}^{3}$ and the hyperbolic $3$-space with negative constant sectional curvature not equal to 1-1. sectional curvature 1-1 and the related volume element $dv$. For a fixed point $x_{0}\in\mathbb{H}^{3}$, set $r:=d_{\mathbb{H}^{3}}(x_{0},y)$, which measures the hyperbolic distance between two points $x_{0}$, $y$ in $\mathbb{H}^{3}$. If the potential function $\phi$ was chosen to be
$\phi(r)=-2\log\left(\frac{r}{\sinh r}\right),$
i.e. $e^{-\phi(r)}=\left(\frac{r}{\sinh r}\right)^{2}$, then the weighted heat kernel is given by
$H^{\phi}(x_{0},y,t)=H^{\phi}(r,t)$ $=$ $\frac{e^{-t/4}}{(4\pi t)^{3/2}}\frac{r}{\sinh r}\times\exp\left(-\frac{r^{2}}{4t}\right)$
$=$ $\frac{e^{-t/4}}{(4\pi t)^{3/2}}\frac{d_{\mathbb{H}^{3}}(x_{0},y)}{\sinh(d_{\mathbb{H}^{3}}(x_{0},y))}\times\exp\left(-\frac{d^{2}_{\mathbb{H}^{3}}(x_{0},y)}{4t}\right).$
Readers can check [38, Subsection 4.6] for the information on the weighted heat kernel $H^{\phi}$ in hyperbolic spaces with other dimensions.
Remark 1.10.
From Example 1.7 and Remark 1.8, one easily knows that for a nontrivial potential function $\phi$, it would also be more complicated if $\phi$ was additionally required to be radial. From this aspect, it should be important and meaningful to get possible comparisons for the weighted heat kernel as what we have done in Theorem 1.6.
Remark 1.11.
In (1) of Theorem 1.6, $\phi$ should be a radial function of the Riemannian distance parameter $s=d_{M^{n}}(q,\cdot)=d_{M^{-}}(q^{-},\cdot)$ with the starting point $q$ on $M^{n}$ or the starting point $q^{-}$ on the model manifold $M^{-}$. However, in (2) of Theorem 1.6, $\phi$ should be a radial function of $s=d_{M^{n}}(q,\cdot)=d_{M^{+}}(q^{+},\cdot)$ with the starting point $q$ on $M^{n}$ or the starting point $q^{+}$ on the model manifold $M^{+}$.
Remark 1.12.
One might see that $M^{+}$ coincides with $M^{-}$ if the (lower or upper) bound function $\kappa(s)$ for the radial Ricci curvature or for the radial sectional curvature was the same, since in this case one can get the same warping function $f(s)$ by solving the initial value problem (1.5).
Remark 1.13.
It should be interesting to know:
  • When does the initial value problem (1.5) have a unique solution on $(0,\infty)$?
That is to say, under what kind of assumptions, one has $l=\infty$. This question is important because if $l=\infty$, the weighted heat kernel comparison (1.7) or (1.8) might sustain continuously without any interruption. By the Sturm-Picone comparison theorem, it is easy to know that if $\kappa(s)\leq 0$, then (1.5) has a unique solution on $(0,\infty)$, i.e. $l=\infty$. What about other situations? In [12], [48, Chapter 2, Section 2.6], some sufficient conditions involving $\kappa(s)$ have been given such that the initial value problem (1.5) has a unique solution on $(0,\infty)$. To the best of our knowledge, so for, the sufficient and necessary condition of the result $l=\infty$ for the initial value problem (1.5) has not been given yet.
Corollary 1.14.
If $\phi=const.$, then the weighted heat kernel $H^{\phi}(\cdot,\cdot,\cdot)$ becomes the usual heat kernel $H(\cdot,\cdot,\cdot)$ of the heat equation, and correspondingly, Theorem 1.6 degenerates into our previous work [49, Theorem 6.6] exactly.
Remark 1.15.
In [49, Theorem 6.6], the author has given lower and upper bounds for the heat kernel of geodesic balls on complete manifolds with radial (Ricci and sectional) curvatures bounded. This conclusion extends Cheeger-Yau’s lower bound estimate [19] and Debiard-Gaveau-Mazet’s upper bound estimate [28], respectively. Hence, one can easily see from Corollary 1.14 that our Theorem 1.6 here improves the heat kernel estimates in [19, 28, 49] a lot.
Remark 1.16.
As mentioned in [49, Remark 6.7], the completeness of the prescribed manifold $M^{n}$ is a little strong to get comparison results for the heat kernel, since Cheeger-Yau [19] have shown that if the injectivity radius at some point $q$ of a prescribed manifold $M^{n}$ is bounded from below, then under the assumptions on curvature therein, a lower bound can be given for the heat kernel of geodesic balls on $M^{n}$. Maybe the completeness of $M^{n}$ is also a little strong to get comparisons (1.7) and (1.8) for the weighted heat kernel. However, here we wish to assume that $M^{n}$ is complete, this is because that if $M^{n}$ is complete, then for $B(q,r_{0})\subseteq M^{n}$ with $r_{0}$ finite we can always find optimally continuous bounds for the radial Ricci and sectional curvatures with respect to $q$ (see (2.14) and (2.15)). Therefore, the assumption on the completeness of $M^{n}$ is feasible.
On $(M^{n},\langle\cdot,\cdot\rangle,e^{-\phi}dv)$, for a bounded domain $\Omega\subset M^{n}$, it is natural to consider the following Dirichlet eigenvalue problem of the Witten-Laplacian
$\left\{\begin{array}[]{ll}\Delta_{\phi}u+\lambda u=0\qquad&\mathrm{in}~~\Omega,\\[1.42262pt] u=0\qquad&\mathrm{on}~~\partial\Omega,\end{array}\right$. (1.11)
with $\partial\Omega$ the boundary777 In general case, to avoid a little bit boring discussion on the regularity of the boundary $\partial\Omega$, which is obviously not the point mainly considered in this paper, one might assume that $\partial\Omega$ is smooth for the Dirichlet eigenvalue problem of the Witten-Laplacian. However, if one checks carefully the details in this paper, he or she would find that weaker regularity assumptions on $\partial\Omega$, such as $\partial\Omega$ is piecewise smooth or Lipschitz continuous, can also make sure that the spectral results of this paper are valid. Based on this reason, in this paper, we do not make any explicit assumption for the regularity of $\partial\Omega$ and we hope that one could understand that a suitable regularity assumption for $\partial\Omega$ has been used potentially in our paper. of the bounded domain $\Omega$. For this problem, one has (see e.g. [21, pp. 1247-1248]):
  • The eigenvalue problem (1.11) only has discrete spectrum, and all the elements (i.e. eigenvalues) in this discrete spectrum can be listed non-decreasingly as follows
    $0<\lambda_{1,\phi}(\Omega)<\lambda_{2,\phi}(\Omega)\leq\lambda_{3,\phi}(\Omega)\leq\cdots\uparrow+\infty.$ (1.12)
    Each eigenvalue888 For convenience and without any confusion, except specification we wish to write $\lambda_{i,\phi}(\Omega)$ as $\lambda_{i,\phi}$ directly. This abbreviation convention would also be used for eigenvalues of other types discussed in the sequel. $\lambda_{i,\phi}$, $i=1,2,\ldots$, in the sequence (1.12) was repeated according to its multiplicity (which is finite and equals the dimension of the eigenspace of $\lambda_{i,\phi}$).
  • By the variational principle, it is not hard to see that the first Dirichlet eigenvalue $\lambda_{1,\phi}(\Omega)$ can be characterized as follows
    $\lambda_{1,\phi}(\Omega)=\inf\left\{\frac{\int_{\Omega}|\nabla f|^{2}e^{-\phi}dv}{\int_{\Omega}f^{2}e^{-\phi}dv}=\frac{\int_{\Omega}|\nabla f|^{2}d\mu}{\int_{\Omega}f^{2}d\mu}\Bigg|f\in W^{1,2}_{0,\phi}(\Omega),f\neq 0\right\}.$ (1.13)
    Using notations as in [21], here $W^{1,2}_{0,\phi}(\Omega)$ denotes a Sobolev space, which is the completion of the set of smooth functions (with compact support) $C^{\infty}_{0}(\Omega)$ under the following Sobolev norm
    $\|f\|^{\phi}_{1,2}:=\left(\int_{\Omega}f^{2}e^{-\phi}dv+\int_{\Omega}|\nabla f|^{2}e^{-\phi}dv\right)^{1/2}.$
    Moreover, the $k$-th Dirichlet eigenvalue $\lambda_{k,\phi}(\Omega)$ can be characterized as follows
    $\lambda_{k,\phi}(\Omega)=\inf\left\{\frac{\int_{\Omega}|\nabla f|^{2}d\mu}{\int_{\Omega}f^{2}d\mu}\Bigg|f\in W^{1,2}_{0,\phi}(\Omega),f\neq 0,\int_{\Omega}ff_{i}d\mu=0\right\},$
    where $f_{i}$, $i=1,2,\cdots,k-1$, denotes an eigenfunction of $\lambda_{i,\phi}(\Omega)$.
In our previous work [22, Introduction], we have given a detailed explanation on why is interesting and important to study spectral geometric problems related to the Witten-Laplacian. We already have some interesting works about spectral isoperimetric inequalities, spectral estimates and geometric functional inequalities related to the Witten-Laplacian – see, e.g., [21, 22, 30, 50, 52, 73].
By the way, if $\phi=const.$, then the eigenvalue problem (1.11) would degenerate into the classical fixed membrane problem of the Laplacian, which has been extensively studied and can be used as a mathematical model to describe the vibration of a membrane with boundary fixed (when $M^{2}=\mathbb{R}^{2}$, and moreover, in this case, $\Omega$ is a bounded planar domain). We wish to refer readers to several books [16, 54, 67] for a relatively comprehensive knowledge about the history and some progresses of spectral problems of the Laplacian.
More than 10 years ago, P. Freitas, J. Mao, and I. Salavessa [33] successfully extended the classical Bishop’s volume comparison theorems to a more general setting (see also Section 2 for details), that is, the curvature assumptions for volume comparisons have been weakened from (Ricci or sectional) curvature bounded (by a constant) to radial (Ricci or sectional) curvature bounded (by some continuous function with respect to the Riemannian distance parameter), and then they improved S. Y. Cheng’s eigenvalue comparison theorems in [24, 25] by applying their volume comparisons and using spherically symmetric manifolds as the model space.
As applications of the weighted heat kernel comparison (i.e. Theorem 1.6), we can get the following eigenvalue comparison theorems of the Witten-Laplacian, which can be seen as the weighted version of those shown in [33, Theorems 3.6 and 4.4].
Theorem 1.17.
Assume that an $n$-dimensional ($n\geq 2$) complete Riemannian manifold $M^{n}$ has a radial Ricci curvature lower bound $(n-1)\kappa(s)$ w.r.t. $q\in M^{n}$, where $s=d_{M^{n}}(q,\cdot)$ denotes the Riemannian distance from $q$. Assume that the potential function $\phi$ has Property 1. Then for $r_{0}<\min\{\ell(q),l\}$ with $\ell(q)$ defined as in (2.7), one has
$\lambda_{1,\phi}(B(q,r_{0}))\leq\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right),$ (1.14)
and the equality holds if and only if $B(q,r_{0})$ is isometric to $\mathscr{B}_{n}(q^{-},r_{0})$.
Theorem 1.18.
Assume that an $n$-dimensional ($n\geq 2$) complete Riemannian manifold $M^{n}$ has a radial sectional curvature upper bound $\kappa(s)$ w.r.t. $q\in M^{n}$, and assume that the potential function $\phi$ has Property 1. Then for $r_{0}<\min\{\mathrm{inj}(q),l\}$ with $\mathrm{inj}(q)$ defined as in (2.5), one has
$\lambda_{1,\phi}(B(q,r_{0}))\geq\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{+},r_{0})\right),$ (1.15)
and the equality holds if and only if $B(q,r_{0})$ is isometric to $\mathscr{B}_{n}(q^{+},r_{0})$.
Remark 1.19.
Denote by $\lambda_{k}(\cdot)$ the $k$-th Dirichlet eigenvalue of the Laplacian on a given geometric object. Clearly, if $\phi=const.$, then (1.14)-(1.15) reduce to
$\lambda_{1}(B(q,r_{0}))\leq\lambda_{1}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$ (1.16)
and
$\lambda_{1}(B(q,r_{0}))\geq\lambda_{1}\left(\mathscr{B}_{n}(q^{+},r_{0})\right),$ (1.17)
respectively. Moreover, when the equalities in the above two eigenvalue inequalities hold, the corresponding isometric rigidity between two geodesic balls can also be attained. This conclusion is exactly the assertions in [33, Theorems 3.6 and 4.4].
Remark 1.20.
In Appendix B, another approach of proving Theorem 1.17 and Theorem 1.18 would be shown carefully. All the calculations in Appendix B were carried out by my master student J. J. Gu under my guidance, and would constitute the main content of her master degree’s dissertation [39] which will be formally issued in June 2026. By the way, the assertion in Theorem 1.17 will be generalized to the case of nonlinear weighted $p$-Laplacian ($1<p<\infty$) in Appendix B as well.
Remark 1.21.
Specially, if curvature assumptions were strengthened to be $\mathrm{Ric}(M^{n})\geq(n-1)K$ or $\mathrm{Sec}(M^{n})\leq K$ for some constant $K$ (i.e. the Ricci (or sectional) curvature of the complete $n$-manifold $M^{n}$ is bounded from below (or above) by some constant), then the eigenvalue comparisons (1.16)-(1.17) would degenerate into
$\lambda_{1}(B(q,r_{0}))\leq\lambda_{1}\left(\mathcal{B}_{n}(K,r_{0})\right)$ (1.18)
and
$\lambda_{1}(B(q,r_{0}))\geq\lambda_{1}\left(\mathcal{B}_{n}(K,r_{0})\right),$ (1.19)
where $\mathcal{B}_{n}(K,r_{0})$ denotes a geodesic ball in the $n$-dimensional space form $\mathbb{M}^{n}(K)$ having constant sectional curvature $K$. Moreover, the equality in (1.18) or (1.19) holds if and only if $B(q,r_{0})$ is isometric to $\mathcal{B}_{n}(K,r_{0})$. Since any space form is two-points homogeneous, there is no need to mention any information for the center of the geodesic ball $\mathcal{B}_{n}(K,r_{0})$. Eigenvalue comparisons (1.18)-(1.19) and related rigidity results are main conclusions in S. Y. Cheng’s works [24, 25]. Based on this reason, we wish to call (1.14)-(1.17) “Cheng-type eigenvalue comparisons”. In fact, this naming convention has already been used in our previous works [33, 48, 49].
Remark 1.22.
In [63, Section 4], by mainly using a similar argument to that in Cheng’s proof of [24, Theorem 1.1], A. G. Setti obtained an eigenvalue comparison theorem for the first Dirichlet eigenvalue of the Witten-Laplacian on geodesic balls of complete Riemannian $n$-manifolds with a modified Ricci curvature bounded from below by a constant. More precisely, for a complete Riemannian $n$-manifold $M^{n}$ with the metric $\langle\cdot,\cdot\rangle$, let $w$ be a given smooth strictly positive function on $M^{n}$ (which was called weight function therein), and then A. G. Setti defined a modified Ricci curvature as follows
$\mathrm{Ric}_{w}:=\mathrm{Ric}-w^{-1}\mathrm{Hess}w.$ (1.20)
Since $w$ is a smooth strictly positive function on $M^{n}$, it is easy to know that there must exist a potential function $\phi$ (which is smooth real-valued) on $M^{n}$ such that $w=e^{-\phi}$, and then one has from (1.20) that
$\mathrm{Ric}_{w}=\mathrm{Ric}-e^{\phi}\mathrm{Hess}(e^{-\phi})=\mathrm{Ric}+\mathrm{Hess}\phi-d\phi\otimes d\phi=\mathrm{Ric}_{\phi}-d\phi\otimes d\phi.$ (1.21)
Therefore, it follows from (1.21) that comparing with the Bakry–Émery Ricci curvature tensor $\mathrm{Ric}_{\phi}$, the modified Ricci curvature tensor $\mathrm{Ric}_{w}$ defined in [63] has an extra term $d\phi\otimes d\phi$. Besides, the only difference between $\mathrm{Ric}_{w}$ and the $N$-dimensional Bakry-Émery Ricci curvature tensor $\mathrm{Ric}_{\phi}^{N}$ is the coefficient $\frac{1}{N-n}$ in front of the term $d\phi\otimes d\phi$, which in my opinion is not an essential difference. In [63, Theorem 4.2], A. G. Setti showed that if
$\mathrm{Ric}_{w}(v_{x},v_{x})\geq\alpha$
for some constant $\alpha$ in the geodesic ball $B(q,R)$ on a given complete Riemannian $n$-manifold $M^{n}$, where $\forall x\in B(q,R)$ and $v_{x}$ is radial unit tangent vector at $x$ defined as (2.6), then
$\lambda_{1,\phi}(B(q,R))\leq\lambda_{1}\left(\mathcal{B}_{n+1}(\alpha/n,R)\right)$ (1.22)
with the weight function $w$ satisfying $w=e^{-\phi}$. That is to say, the first Dirichlet eigenvalue $\lambda_{1,\phi}(B(q,R))$ of the Witten-Laplacian $\Delta_{\phi}$ on the geodesic ball $B(q,R)\subseteq M^{n}$ is bounded from above by the first Dirichlet eigenvalue $\lambda_{1}\left(\mathcal{B}_{n+1}(\alpha/n,R)\right)$ of the Laplacian on a geodesic ball $\mathcal{B}_{n+1}(\alpha/n,R)$ in the $(n+1)$-dimensional space form $\mathbb{M}^{n+1}(\alpha/n)$. Specially, in [63, Section 4], he also mentioned that if furthermore $\alpha>0$, then using a Bishop-type volume comparison theorem [63, Theorem 4.1], one knows that $d_{M^{n}}(q,\hat{q})\leq\pi\sqrt{\frac{n}{\alpha}}$, $\forall\hat{q}\in M^{n}$. Hence, if $R>\pi\sqrt{\frac{n}{\alpha}}$, then $B(q,R)=M^{n}$ and $\mathcal{B}_{n+1}(\alpha/n,R)=\mathbb{M}^{n+1}(\alpha/n)$, which implies $\lambda_{1,\phi}(B(q,R))=0=\lambda_{1}\left(\mathcal{B}_{n+1}(\alpha/n,R)\right)$, and the comparison (1.22) holds trivially. That is to say, the eigenvalue comparison (1.22) does not have the rigidity characterization for the equality “==”, however in some special cases the equality can be achieved.
Recently, for a complete Riemannian $n$-manifold $M^{n}$ and a point $q\in M^{n}$, if the modified Ricci curvature $\mathrm{Ric}_{w}$ satisfies
$\mathrm{Ric}_{w}(v_{x},v_{x})\geq\kappa(r(x)),$
where $\forall x\in B(q,R)$, $r(x)=d_{M^{n}}(q,x)$, and $\kappa:(0,R)\rightarrow\mathbb{R}$ is a continuous function of the radial distance parameter, then we [20] can improve Setti’s eigenvalue comparison (1.22) to a more general setting, and the corresponding model space becomes $(n+1)$-dimensional spherically symmetric manifolds.
Remark 1.23.
The reason we do not put also the eigenvalue comparison results [20] in this paper is that (1.14)-(1.15) are the comparisons between eigenvalues of the same type (i.e. the first Dirichlet eigenvalues of the Witten-Laplacian on different geodesic balls), while the ones in [20] are comparisons between Dirichlet eigenvalues of different elliptic differential operators.
Remark 1.24.
The study of extremum problems is one of core topics in Spectral Geometry, and we wish to refer readers to A. Henrot’s book [41] for a relatively comprehensive knowledge about this topic. As briefly mentioned before, for an $n$-dimensional($n\geq 2$) Riemannian manifold $M^{n}$ and a bounded domain $\Omega$ on $M^{n}$, if the potential function $\phi$ was assumed to be radial, suitably concave or convex, then recently R. F. Chen and J. Mao [21] successfully and completely solved the extremum problem
$\min\{\lambda_{k,\phi}(\Omega)|\Omega\subset M^{n},|\Omega|_{n,\phi}=const.\}$
when $k=1,2$ and $M^{n}\equiv\mathbb{M}^{n}(K)$, i.e. the $n$-dimensional space form with constant sectional curvature $K$. This conclusion definitely extends those classical spectral isoperimetric inequalities, i.e. the Faber-Krahn inequality, the Hong-Krahn-Szegő inequality of the Laplacian under the constraint of fixed volume. From the viewpoint of extremum problems, our Theorems 1.17-1.18 here give characterizations for the extremum value of the functional
$B(q,r_{0})\mapsto\lambda_{1,\phi}(B(q,r_{0}))$
provided the potential function $\phi$ satisfying Property 1, where $B(q,r_{0})$ is the geodesic ball with center $q$ and fixed radius $r_{0}$ on a given complete Riemannian $n$-manifold. Speaking in other words, if the potential function $\phi$ satisfies Property 1, then the value of the first Dirichlet eigenvalue $\lambda_{1,\phi}(B(q,r_{0}))$ of the Witten-Laplacian would change as long as the radial (Ricci or sectional) curvature with respect to $q$ changes.
Comparing with our research experience in [21], it is so surprising that except the radial curvature assumptions, we can get the Cheng-type eigenvalue comparisons in Theorems 1.17-1.18 only assuming that the potential function $\phi$ satisfies Property 1.
Remark 1.25.
Specially, when $n=2$, $M^{2}$ is a complete surface, and as mentioned in Section 2 below, for any $q\in M^{2}$, the radial Ricci and sectional curvatures become exactly the Gaussian curvature, and one can always find sharp lower and upper bounds $\kappa_{\pm}(q,s)$ for the curvature assumptions with respect to the given point $q$, and then it follows from Theorems 1.17-1.18 that the functional
$B(q,s)\mapsto\lambda_{1,\phi}(B(q,s))$
would attain its minimal value if the Gaussian curvature in the geodesic disk $B(q,s)$ vanishes identically, where $s\in(0,\infty)$ and the potential function $\phi$ satisfies Property 1.
This paper is organized as follows. Some preliminaries and useful facts will be introduced in Section 2. A proof of Theorem 1.6 (i.e. a weighted heat kernel comparison theorem for the weighted heat equation on complete manifolds with radial curvatures bounded), together with some basic properties related to the weighted heat kernel, will be given in Section 3. As an application, in Section 4 we will mainly use Theorem 1.6 to give a proof for two Cheng-type eigenvalue comparisons (1.14) in Theorem 1.17 and (1.15) in Theorem 1.18. The detailed calculations for the weighted heat kernel given in Example 1.7 will be given in Appendix A. We shall give anther proof for eigenvalue comparisons (1.14)-(1.15) and their rigidity characterizations (i.e. Theorems 1.17 and 1.18) in Appendix B.
On $1^{\mathrm{st}}$ March 2026, we posted the first version of this paper on arXiv. However, on $12^{\mathrm{th}}$ March 2026, we noticed the work [64] which was just posted on arXiv on $9^{\mathrm{th}}$ March 2026 (i.e. almost three days ago), and has cited my two works [29, 49]. We found that the assertion in Theorem 1.18 can be improved to the case of weighted $p$-Laplacian by making some adjustments to the argument in the proof of [64, Theorem 1.4]. We wish to write this content as an Appendix C in the second version of this paper.
2 Preliminaries and some useful facts
Given an $n$-dimensional ($n\geq 2$) complete Riemannian manifold $M^{n}$ with the metric $\langle\cdot,\cdot\rangle$, for any point $q\in M^{n}$, the exponential map $\exp_{q}:\mathcal{D}_{p}\to M^{n}\setminus Cut(q)$ gives a diffeomorphism from a star-shaped open set $\mathcal{D}_{p}$ of the tangent space $T_{q}M^{n}$ with
$\mathcal{D}_{q}=\left\{s\xi|~0\leq{s}<d_{\xi},~\xi\in{S^{n-1}_{q}}\right\}$
onto the open set $M^{n}\setminus Cut(q)$, where $Cut(q)$ is the cut locus of $q$, $S^{n-1}_{q}$ denotes the unit sphere of $T_{q}M^{n}$, and $d_{\xi}$ is defined by
$d_{\xi}=d_{\xi}(q):=\sup\{r>0|~\gamma_{\xi}(s)=\gamma_{(q,\xi)}(s):=\exp_{q}(s\xi)~{\rm{is~the~unique}}$
$\qquad\qquad\qquad\qquad\qquad\qquad{\rm{minimal~geodesic~joining}}~q~{\rm{and}}~\gamma_{\xi}(r)\}.$ (2.1)
For a fixed unit vector $\xi\in S^{n-1}_{q}\subset T_{q}M^{n}$, let $\xi^{\bot}$ be the orthogonal complement of $\{\mathbb{R}\xi\}$ in $T_{q}M^{n}$ and $\tau_{s}:T_{q}M^{n}\rightarrow{T_{\exp_{q}(s\xi)}M^{n}}$ be the parallel translation along $\gamma_{\xi}$. The path of linear transformations $\mathbb{A}(s,\xi):\xi^{\bot}\rightarrow{\xi^{\bot}}$ is given by
$\mathbb{A}(s,\xi)\eta=(\tau_{s})^{-1}Y_{\eta}(s),$
where $Y_{\eta}(s)=d(\exp_{q})_{(s\xi)}(s\eta)$ is the Jacobi field along $\gamma_{\xi}$ satisfying $Y_{\eta}(0)=0$, and $(\nabla_{s}Y_{\eta})(0)=\eta$. This operator satisfies the Jacobi equation $\mathbb{A}^{\prime\prime}+\mathcal{R}\mathbb{A}=0$ with initial conditions $\mathbb{A}(0,\xi)=0$, $\mathbb{A}^{\prime}(0,\xi)=I$, where $\mathcal{R}(s)$ is the self-adjoint operator on $\xi^{\bot}$, $\mathcal{R}(s)\eta=(\tau_{s})^{-1}R(\gamma^{\prime}_{\xi}(s),\tau_{s}\eta)\gamma^{\prime}_{\xi}(s)$ with $R$ the curvature tensor on $M^{n}$. The trace of the later operator is just the radial Ricci tensor
$\mathrm{Ric}_{\gamma_{\xi}(s)}(\gamma^{\prime}_{\xi}(s),\gamma^{\prime}_{\xi}(s))$
along unit speed geodesics starting from $q$. Gauss’s lemma tells us that the Riemannian metric on $M^{n}\setminus Cut(q)$ in the spherical geodesic coordinate chart can be expressed as
$g:=\langle\cdot,\cdot\rangle_{\exp_{q}(s\xi)}=ds^{2}+|\mathbb{A}(s,\xi)d\xi|^{2},\quad\forall~s\xi\in\mathcal{D}_{q}.$ (2.2)
We consider the metric components $g_{ij}(s,\xi)$, $i,j\geq 1$, in a coordinate system $\{s,\xi_{a}\}_{a\geq 2}$ formed by fixing an orthonormal basis $\{\eta_{a}\}_{a\geq 2}$ of $\xi^{\bot}=T_{\xi}S^{n-1}_{q}$, and extending it to a local frame $\{\xi_{a}\}_{a\geq 2}$ of $S_{q}^{n-1}$. Define a function $J>0$ on $\mathcal{D}_{q}$ by
$J^{n-1}=\sqrt{|g|}:=\sqrt{\det[g_{ij}]},$ (2.3)
that is, $\sqrt{|g|}=\det\mathbb{A}(t,\xi)$, and $dv=J^{n-1}dsd\sigma$ is the volume element of $M^{n}\setminus Cut(q)$, where $d\sigma$ denotes the (n1)(n-1)-dimensional volume element on $\mathbb{S}^{n-1}\equiv S_{q}^{n-1}\subseteq{T_{q}M^{n}}$. Since $\tau_{s}:T_{q}M^{n}\rightarrow{T_{\exp_{q}(s\xi)}M^{n}}$ is an isometry, one has
$\langle d(\exp_{q})_{(s\xi)}(s\eta_{a}),d(\exp_{q})_{(s\xi)}(s\eta_{b})\rangle=\langle\mathbb{A}(s,\xi)(\eta_{a}),\mathbb{A}(s,\xi)(\eta_{b})\rangle,$
and so $\sqrt{|g|}=\det\mathbb{A}(s,\xi)$. Therefore, by applying (2.2) and (2.3), it follows that the volume $\mathrm{vol}(B(q,r))$ of the geodesic ball $B(q,r)$ on $M^{n}$ is given by
$\mathrm{vol}(B(q,r))=\int_{S_{q}^{n-1}}\int_{0}^{\min\{r,d_{\xi}\}}\sqrt{|g|}dsd\sigma=\int_{S_{q}^{n-1}}\left(\int_{0}^{\min\{r,d_{\xi}\}}\det(\mathbb{A}(s,\xi))ds\right)d\sigma.$ (2.4)
Let
$\mathrm{inj}(q):=d_{M^{n}}(q,Cut(q))=\min\limits_{\xi\in S_{q}M^{n}}d_{\xi}$ (2.5)
be the injectivity radius at $q$. Generally, one has $B(q,\mathrm{inj}(q))\subseteq M^{n}\setminus\mathrm{Cut}(q)$. Besides, for $r<\mathrm{inj}(q)$, one has
$\mathrm{vol}(B(q,r))=\int_{0}^{r}\int_{S_{q}^{n-1}}\det(\mathbb{A}(s,\xi))d\sigma ds.$
If $r(x)=d_{M^{n}}(x,q)$ denotes the intrinsic distance to the point $q$, $x\in M^{n}\setminus(Cut(q)\cup\{q\})$, then the unit vector field
$v_{x}=\nabla r(x)$ (2.6)
is the radial unit tangent vector at $x$, according to the definition given in [42]. This is because for any $\xi\in S_{q}^{n-1}$ and $s_{0}>0$, one has $\nabla r(\gamma_{\xi}(s_{0}))=\gamma^{\prime}_{\xi}(s_{0})$ when the point $\gamma_{\xi}(s_{0})=\exp_{q}(s_{0}\xi)$ is away from the cut locus of $q$ (see e.g. [37]). Set
$\ell(q):=\sup\limits_{x\in M^{n}}r(x).$ (2.7)
One has $\ell(q)=\max_{\xi\in S_{q}M^{n}}d_{\xi}$ (see page 704 of [33, Section 2] for the explanation). Clearly,
$\ell(q)\geq\mathrm{inj}(q).$
We also recall the following inequality about $r(x)$ (cf. [61], Prop. 39, and pp. 266-267), with $\partial_{r}=\nabla r$ as a vector of differentiation (see Prop. 7 on p. 47 of the same reference),
$\partial_{r}\Delta r\leq\partial_{r}\Delta r+\|\mathrm{Hess}\,r\|^{2}=-\mathrm{Ric}(\partial_{r},\partial_{r}),\quad\quad\mbox{with}\quad\Delta r=\partial_{r}\ln(\sqrt{|g|}),$
which implies that
$J^{\prime\prime}+\frac{1}{(n-1)}\,\mathrm{Ric}{(\gamma_{\xi}^{\prime}(t),\gamma_{\xi}^{\prime}(t))}\,J\leq 0,$ (2.8)
$J(0,\xi)=0,\quad J^{\prime}(0,\xi)=1.$ (2.9)
(2.8) and (2.9) make a fundamental role in the derivation of the so-called generalized Bishop’s volume comparison theorem I (see Theorem 2.6 below), which has been proven in our previous works [33, 48].
We need the following concept:
Definition 2.1.
A domain $\Omega=\exp_{q}([0,l)\times{S}_{q}^{n-1})\subset M^{n}\backslash Cut(q)$, with $l<inj(p)$, is said to be spherically symmetric with respect to a point $q\in\Omega$, if and only if the matrix $\mathbb{A}(s,\xi)$ satisfies $\mathbb{A}(s,\xi)=f(s)I$, for a function $f\in{C^{2}([0,l))}$, with $f(0)=0$, $f^{\prime}(0)=1$, and $f|_{(0,l)}>0$.
Obviously, on the domain $\Omega$ given as in Definition 2.1 the Riemannian metric of $M^{n}$ can be expressed by
$\langle\cdot,\cdot\rangle=ds^{2}+f^{2}(s)|d\xi|^{2},\qquad\forall~\xi\in S_{q}M^{n},~0\leq s<l,$ (2.10)
[…]
$(s,\xi)\thicksim(t,\eta)\Longleftrightarrow\left\{\begin{array}[]{ll}s=t~\mathrm{and}~\xi=\eta,~~~\mathrm{or}\\ s=t=0.\end{array}\right$.
This equivalent relation is natural, and usually one can just use $0,l)\times_{f(s)}\mathbb{S}^{n-1}$ to represent this quotient space. That is to say, $M^{\ast}=[0,l)\times_{f(s)}\mathbb{S}^{n-1}$ with $f$ satisfying conditions in Definition 2.1 is a spherically symmetric manifold with $q=\{0\}\times_{f}\mathbb{S}^{n-1}$ the base point and (2.10) as its metric. For $M^{\ast}$ and $r<l$, by (2.4) it follows that the volume of the geodesic ball $\mathscr{B}(q,r)\subset M^{\ast}$ is given by
$\mathrm{vol}(\mathscr{B}(q,r))=w_{n}\int_{0}^{r}f^{n-1}(s)ds,$
and moreover, the volume of the boundary $\partial\mathscr{B}(q,r)$ is given by
$\mathrm{vol}(\partial\mathscr{B}(q,r))=w_{n}f^{n-1}(r),$
where $w_{n}$ denotes the (n1)(n-1)-volume of the unit sphere $\mathbb{S}^{n-1}$.
We also need the following concepts, which can be found e.g. in our previous works [33, 48, 49].
Definition 2.2.
Given a continuous function $\kappa:[0,l)\rightarrow\mathbb{R}$, we say that $M^{n}$ has a radial Ricci curvature lower bound $(n-1)\kappa$ at the point $q$ if
$\mathrm{Ric}(v_{x},v_{x})\geq(n-1)\kappa(r(x)),\quad\forall x\in M^{n}\backslash Cut(q)\cup\{q\},$ (2.12)
where $\mathrm{Ric}$ is the Ricci curvature of $M^{n}$.
Definition 2.3.
Given a continuous function $\kappa:[0,l)\rightarrow\mathbb{R}$, we say that $M^{n}$ has a radial sectional curvature upper bound $\kappa$ at the point $q$ if
$\mathscr{K}(v_{x},V)\leq\kappa(r(x)),\quad\forall x\in M^{n}\setminus Cut(q)\cup\{q\},$ (2.13)
where $V\perp{v_{x}}$, $V\in{S^{n-1}_{x}}\subseteq{T_{x}M^{n}}$, and $\mathscr{K}(v_{x},V)$ denotes the sectional curvature of the plane spanned by $v_{x}$ and $V$.
Remark 2.4.
As pointed out in [33, Remark 2.4] or [49, Remark 2.5], since the radial distance is $r(x)=d_{M^{n}}(x,q)$ for $x=\gamma_{\xi}(s)=\exp_{q}(s\xi)$, the parameter $s$ may be seen as the argument of the continuous function $\kappa:[0,l)\rightarrow\mathbb{R}$ in Definition 2.2 or Definition 2.3. Therefore, one has $\frac{d}{ds}|_{x}=\nabla r(x)=v_{x}$, which implies (2.12) and (2.13) become
$\mathrm{Ric}(\frac{d}{ds},\frac{d}{ds})\geq(n-1)\kappa(s)$
and
$\mathscr{K}(\frac{d}{ds},V)\leq\kappa(s),$
respectively. By the way, since $x\in M^{n}\backslash Cut(q)\cup\{q\}$, one knows that the geodesic $\gamma_{\xi}(s)=\exp_{q}(s\xi)$, emanating from $q$ and joining $x$, should be unit-speed and minimizing. So, if a given manifold $M^{n}$ satisfies (2.12) (resp., (2.13)), then we say that $M^{n}$ has a radial Ricci curvature lower bound $(n-1)\kappa$ (resp., a radial sectional curvature upper bound $\kappa$) along any unit-speed minimizing geodesic starting from the point $q\in M^{n}$, that is to say, for convenience and simplification, its radial Ricci curvature is bounded from below with respect to $q$ (resp., radial sectional curvature is bounded from above with respect to $q$) by the continuous function $(n-1)\kappa$ (resp., $\kappa$).
Remark 2.5.
Similarly, one can define the concept that the radial Ricci/sectional curvature has an upper/lower bound (with respect to some point) which is given by a continuous function of the radial distance parameter. By the way, the concept that the radial (Ricci or sectional) curvature is bounded at some point has been improved to a more general setting, i.e. the integral radial (Ricci or sectional) curvature is bounded with respect to some point – see [53, Definitions 2.1 and 2.2]. In fact, the quantities defined in [53, Definitions 2.1 and 2.2] can be used to measure the deviation of the radial (Ricci or sectional) curvature (with respect to some point $q\in M^{n}$) from a prescribed continuous function $(n-1)\kappa(r(x))$ or $\kappa(r(x))$, where $r(x)=d_{M^{n}}(x,q)$ denotes the intrinsic distance to $q$ on the Riemannian $n$-manifold $M^{n}$.
Now, we wish to point out an important truth, that is, for a prescribed $n$-dimensional complete manifold $M^{n}$, one can always construct the optimal continuous functions $\kappa_{\pm}(q,s)$ with respect to a point $q\in M^{n}$, satisfying Definitions 2.2 and 2.3 respectively. This truth has already been shown in our previous works [33, p. 706], [49, pp. 378-379]. However, for convenience to readers, we wish to repeat the construction here. Recall that for $\xi\in S_{q}^{n-1}\subseteq T_{q}M^{n}$, $\gamma_{\xi}(s)=\exp_{q}(s\xi)$ and its derivative $\gamma^{\prime}_{\xi}(s)$ depend smoothly on the variables $(s,\xi)$. Set $\mathbb{D}_{q}:=\{(s,\xi)\in[0,\infty)\times S_{q}^{n-1}|0\leq s<d_{\xi}\}$ with closure $\overline{\mathbb{D}}_{q}:=\{(s,\xi)\in[0,\infty)\times S_{q}^{n-1}|0\leq s\leq d_{\xi}\}$. Then we can define
$\kappa_{-}(q,s):=\min\limits_{\{\xi|(s,\xi\in\overline{\mathbb{D}}_{q})\}}\frac{\mathrm{Ric}_{\gamma_{\xi}(s)}(\gamma^{\prime}_{\xi}(s),\gamma^{\prime}_{\xi}(s))}{n-1},\qquad 0\leq s<\ell(q)$ (2.14)
and
$\kappa_{+}(q,s):=\max\limits_{\{(\xi,V)|\gamma^{\prime}_{\xi}(s)\perp V,|V|=1\}}\mathscr{K}_{\gamma_{\xi}(s)}(\gamma^{\prime}_{\xi}(s),V),\qquad 0\leq s<\mathrm{inj}(q).$ (2.15)
If $\ell(q)<+\infty$, the above two functions can be continuously extended to $s=\ell(q)$ and $s=\mathrm{inj}(q)$ respectively. Furthermore, if $M^{n}$ is closed, then the injectivity radius $\mathrm{inj}(M^{n})=\min_{q\in M^{n}}\mathrm{inj}(q)$ is a positive constant, and then in this case by applying the uniform continuity of continuous functions on compact sets, one knows that $\kappa_{\pm}(q,s)$ are continuous. Hence, for a bounded domain $\Omega\subseteq M^{n}$, one can always find optimally continuous bounds $\kappa_{\pm}(q,s)$ for the radial sectional and Ricci curvatures with respect to the given point $q\in\Omega$. This implies that the assumptions on curvatures in Definitions 2.2 and 2.3 are natural and feasible. Specially, when $n=2$, $M^{2}$ is a complete surface, and then $\kappa_{\pm}(q,s)$ defined by (2.14) and (2.15) are actually the minimum and maximum of the Gaussian curvature on geodesic circles centered at $q$ of radius $s$ on $M^{2}$.
As explained in [33, pp. 706-707] or [49, p. 379], by using the properties of warped product manifolds, one knows that the radial sectional curvature, and the radial component of the Ricci tensor of the spherically symmetric manifold $0,l)\times_{f(s)}\mathbb{S}^{n-1}$ with the base point $q$ are given by
$\mathscr{K}\left(V,\frac{d}{ds}|_{\exp_{q}(s\xi)}\right)=R\left(\frac{d}{ds}|_{\exp_{q}(s\xi)},V,\frac{d}{ds}|_{\exp_{q}(s\xi)},V\right)=-\frac{f^{\prime\prime}(s)}{f(s)}$
for $V\in T_{\xi}\mathbb{S}^{n-1}$, $|V|=1$, and
$\mathrm{Ric}\left(\frac{d}{ds}|_{\exp_{q}(s\xi)},\frac{d}{ds}|_{\exp_{q}(s\xi)}\right)=-(n-1)\frac{f^{\prime\prime}(s)}{f(s)}.$
So, in this setting, Definition 2.2 (resp., Definition 2.3) is satisfied with equality in (2.12) (resp., (2.13)) and correspondingly $\kappa(s)=-f^{\prime\prime}(s)/f(s)$. From the above two facts, we know that in order to define curvature tensor away from $q$, we need to require $f\in C^{2}((0,l))$. Furthermore, if $f^{\prime\prime}(0)=0$ and $f$ is $C^{3}$ at $s=0$, then one has $\lim_{s\rightarrow 0}\kappa(s)=-f^{\prime\prime\prime}(0)$. Although $\nabla r(x)$ is not defined at $x=q$, $\kappa(s)$ is usually required to be continuous at $s=0$, which is equal to require $f$ to be $C^{3}$ at $s=0$.
Now, we wish to recall two Bishop-type volume comparison theorems, which correspond to Theorem 3.3, Corollary 3.4, and Theorem 4.2 in [33] (equivalently, Theorem 2.2.3, Corollary 2.2.4 and Theorem 2.3.2 in [48]). Define a quantity on $M^{n}\setminus Cut(q)$ by
$\theta(s,\xi)=\left[\frac{J(s,\xi)}{f(s)}\right]^{n-1}.$
Let $a(\xi):=\min\{d_{\xi},r_{0}\}$ on $M^{n}$, and we always assume $r_{0}<\min\{\ell(q),l\}$. One has the following facts:
Theorem 2.6 (Generalized Bishop’s comparison theorem I).
(see [33, 48]) Given $\xi\in S_{q}^{n-1}$, and a model space $M^{-}=[0,l)\times_{f}\mathbb{S}^{n-1}$, under the curvature assumption on the radial Ricci tensor, $\mathrm{Ric}(v_{x},v_{x})\geq-(n-1)f^{\prime\prime}(s)/f(s)$, for $x=\gamma_{\xi}(s)=\exp_{q}(s\xi)$ with $s<\min\{d_{\xi},l\}$ (resp. with $s<\min\{a_{\xi},l\}$) the function $\theta(s,\xi)$ is nonincreasing in $s$. In particular, for all $s<\min\{d_{\xi},l\}$  (resp. $s<\min\{a(\xi),l\}$) we have $J(s,\xi)\leq f(s)$. Furthermore, this inequality is strict for all $s\in(s_{0},s_{1}]$, with $0\leq s_{0}<s_{1}<\min\{d_{\xi},l\}$, if the above curvature assumption holds with a strict inequality for $s$ in the same interval. Besides, we have
$\mathrm{vol}(B(q,r_{0}))\leq\mathrm{vol}(\mathscr{B}_{n}(q^{-},r_{0})),$
with equality if and only if $B(q,r_{0})$ is isometric to $\mathscr{B}_{n}(q^{-},r_{0})$.
Theorem 2.7 (Generalized Bishop’s comparison theorem II).
(see [33, 48])  Suppose $M^{n}$ has a radial sectional curvature upper bound given by  $\kappa(s)=-\frac{f^{\prime\prime}(s)}{f(s)}$ for  $s<\beta\leq\min\{inj_{c}(q),l\}$, where $inj_{c}(q)=\inf_{\xi}c_{\xi}$, with $\gamma_{\xi}(c_{\xi})$ a first conjugate point along the geodesic $\gamma_{\xi}(s)=\exp_{q}(s\xi)$. Then on $(0,\beta)$
$\left[\frac{\sqrt{|g|}}{f^{n-1}}\right]^{\prime}\geq 0,\quad\quad\sqrt{|g|}(s)\geq{f^{n-1}(s)},$
and equality occurs in the first inequality at $s_{0}\in(0,\beta)$ if and only if
$\mathcal{R}=-\frac{f^{\prime\prime}(s)}{f(s)},\quad\mathbb{A}=f(s)I$
on all of $0,s_{0}]$.
Remark 2.8.
For any measurable subset $\Gamma$ of $S_{q}M^{n}\subset T_{q}M^{n}$, an annulus $A^{\Gamma}_{r,R}(q)$ can be well-defined as follows
$A^{\Gamma}_{r,R}(q):=\{x\in M^{n}|r\leq r(x)\leq R,$
$\qquad\qquad\qquad{\rm{and~any~minimal~geodesic~\gamma~from}}~q~{\rm{to}}~x~{\rm{satisfies}}~\gamma^{\prime}(0)\in\Gamma\}.$
Clearly, if $r=0$ and $R=r_{0}$, then the annulus $A^{\Gamma}_{0,r_{0}}(q)$ would be exactly the geodesic ball $B(q,r_{0})$ on the given Riemannian $n$-manifold $M^{n}$. In [51, Theorem 3.2] (resp., [51, Theorem 4.3]), we have successfully obtained a relative volume comparison for the annulus $A^{\Gamma}_{r,R}(q)$ on a complete Riemannian $n$-manifold $M^{n}$ having a radial Ricci curvature lower bound $(n-1)f^{\prime\prime}(s)/f(s)$ with respect to the point $q\in M^{n}$ (resp., a radial sectional curvature upper bound $f^{\prime\prime}(s)/f(s)$ with respect to the point $q$). Moreover, those two volume comparison results would immediately imply our Bishop-type volume comparisons (i.e. Theorems 2.6 and 2.7 here) by suitably choosing radius parameters for annuluses – see [51, Corollaries 3.4 and 4.4] for details. In 2023, we attempted to improve our previous volume comparison results mentioned here to the setting that the given complete Riemannian $n$-manifold has integral radial curvature bounds (i.e. a more weaker curvature assumption), and this attempt was successful – see [53, Section 3] for details.
As mentioned in Remark 1.20, the assertion in Theorem 1.17 will be generalized to the case of nonlinear weighted $p$-Laplacian ($1<p<\infty$) in Appendix B. For convenience to readers, here we wish to give some fundamental facts on the Dirichlet eigenvalue problem of the weighted $p$-Laplacian. In fact, for a bounded domain $\Omega$ (with boundary $\partial\Omega$) on the given Riemannian $n$-manifold $M^{n}$, one can consider the following eigenvalue problem
$\left\{\begin{array}[]{ll}\Delta_{p,\phi}u+\lambda|u|^{p-2}u=0\qquad&\mathrm{in}~~\Omega,\\[1.42262pt] u=0\qquad&\mathrm{on}~~\partial\Omega,\end{array}\right$. (2.18)
where $\Delta_{p,\phi}:=e^{\phi}\mathrm{div}(e^{-\phi}|\nabla\cdot|^{p-2}\nabla\cdot)$, $1<p<\infty$, denotes the weighted $p$-Laplacian. That is to say, in a local coordinates chart $\{x^{1},x^{2},\cdots,x^{n}\}$ of $M^{n}$, one has
$\Delta_{p,\phi}u=\frac{e^{\phi}}{\sqrt{\det[g_{ij}]}}\sum\limits_{i,j=1}^{n}\frac{\partial}{\partial x^{i}}\left(e^{-\phi}\sqrt{\det[g_{ij}]}g^{ij}|\nabla u|^{p-2}\frac{\partial u}{\partial x^{j}}\right),$
where $|\nabla u|^{2}=\sum_{i,j=1}^{n}g^{ij}\frac{\partial}{\partial x^{i}}\frac{\partial}{\partial x^{j}}$, and $(g^{ij})=[g_{ij}]^{-1}$ is the inverse of the metric matrix. Clearly, if $\phi=const.$, then $\Delta_{p,\phi}$ would reduce to the classical nonlinear $p$-Laplacian $\Delta_{p}:=\mathrm{div}(|\nabla\cdot|^{p-2}\nabla\cdot)$, and correspondingly, one has
$\Delta_{p}u=\frac{1}{\sqrt{\det[g_{ij}]}}\sum\limits_{i,j=1}^{n}\frac{\partial}{\partial x^{i}}\left(\sqrt{\det[g_{ij}]}g^{ij}|\nabla u|^{p-2}\frac{\partial u}{\partial x^{j}}\right).$
Besides, if $p=2$, then the weighted $p$-Laplacian $\Delta_{p,\phi}$ would become the Witten-Laplacian $\Delta_{\phi}$ exactly. For the eigenvalue problem (2.18), by making suitable adjustments to the arguments in [2, 3, 9, 35, 45] and using the variational principle, it is not hard to get:
  • (2.18) has a positive weak solution, which is unique modulo the scaling, in the space $W^{1,p}_{0,\phi}(\Omega)$, the completion of the set $C^{\infty}_{0}(\Omega)$ of smooth functions compactly supported on $\Omega$ under the Sobolev norm
    $\|u\|^{\phi}_{1,p}:=\left(\int_{\Omega}|u|^{p}e^{-\phi}dv+\int_{\Omega}|\nabla u|^{p}e^{-\phi}dv\right)^{1/p}.$
  • By applying the Ljusternik-Schnirelman principle, there exists a nondecreasing sequence of nonnegative eigenvalues $\{\lambda_{i,p}^{\phi}(\Omega)\}$ tending to $\infty$ as $i\rightarrow\infty$.
  • The first Dirichlet eigenvalue $\lambda_{1,p}^{\phi}(\Omega)$ is simple, isolated, and eigenfunctions associated with $\lambda_{1,p}^{\phi}(\Omega)$ do not change sign. Besides, $\lambda_{1,p}^{\phi}(\Omega)$ can be characterized by
    $\lambda_{1,p}^{\phi}(\Omega)=\inf\left\{\frac{\int_{\Omega}|u|^{p}e^{-\phi}dv}{\int_{\Omega}|\nabla u|^{p}e^{-\phi}dv}=\frac{\int_{\Omega}|u|^{p}d\mu}{\int_{\Omega}|\nabla u|^{p}d\mu}\Big|u\neq 0,~u\in W^{1,p}_{0,\phi}(\Omega)\right\}.$ (2.19)
[…]
In [31, Section 3], we have given several lower bounds for the first eigenvalue of weighted $p$-Laplacian on submanifolds with locally bounded weighted mean curvature.
3 Weighted heat kernel comparison theorems
Given an $n$-dimensional smooth metric measure space $(M^{n},\langle\cdot,\cdot\rangle,e^{-\phi}dv)$ and any bounded domain $\Omega$ in this space, by using the divergence theorem, it is easy to get the following Green’s formula
$\int_{\Omega}h_{1}\Delta_{\phi}h_{2}e^{-\phi}dv=-\int_{\Omega}\langle\nabla h_{1},\nabla h_{2}\rangle e^{-\phi}dv=\int_{\Omega}h_{2}\Delta_{\phi}h_{1}e^{-\phi}dv$
for $h_{1},h_{2}\in C^{\infty}_{0}(\Omega)$, i.e.
$\int_{\Omega}h_{1}\Delta_{\phi}h_{2}d\mu=-\int_{\Omega}\langle\nabla h_{1},\nabla h_{2}\rangle d\mu=\int_{\Omega}h_{2}\Delta_{\phi}h_{1}d\mu.$ (3.1)
It is not hard to check that Green’s formula (3.1) remains valid if the compact support condition for $h_{1},h_{2}$ was replaced by the Neumann boundary condition on the boundary $\partial\Omega$. By (3.1), one knows that the operator $\Delta_{\phi}$ in999 In the weighted setting, similar as the symbol $L^{1}_{\phi}$ in Section 1, $L^{2}_{\phi}$ means doing $L^{2}$ integrals with respect to the weighted measure $d\mu=e^{-\phi}dv$. We will not mention this explanation in the sequel any more. $L^{2}_{\phi}$ with the space $C^{\infty}_{0}(\Omega)$ is symmetric and nonpositive definite. It is natural to ask:
  • Whether the operator $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$ has a self-adjoint extension in $L^{2}_{\phi}$ and whether it is essentially self-adjoint?
This question is crucial for the eigenvalue problem (1.11) and the answer is affirmative. In fact, as mentioned in A. Grigor’yan’s survey [38, Section 2], by suitably making adjustments to those related arguments in e.g. [19, 27, 34, 65], it is not hard to get the following result.
Theorem 3.1.
([38, Section 2]) The operator $\Delta_{\phi}|_{W^{2,2}_{0,\phi}(\Omega)}$ is a self-adjoint nonpositive definite operator in $L^{2}_{\phi}$, where the space $W^{2,2}_{0,\phi}(\Omega)$ is defined as $W^{2,2}_{0,\phi}(\Omega):=\{\rho\in W^{1,2}_{0,\phi}(\Omega)|\Delta_{\phi}\rho\in L^{2}_{\phi}(\Omega)\}$. Moreover, the operator $\Delta_{\phi}|_{W^{2,2}_{0,\phi}(\Omega)}$ is a unique self-adjoint extension of $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$ with the domain in $W^{1,2}_{0,\phi}(\Omega)$. If in addition $\Omega$ is geodesically complete, then the operator $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$ is essentially self-adjoint, i.e., $\Delta_{\phi}|_{W^{2,2}_{0,\phi}(\Omega)}$ is a unique self-adjoint extension of $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$.
Once we have Theorem 3.1 for the extension of the operator $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$. The spectrum of the eigenvalue problem (1.11) is actually the spectral structure of the self-adjoint operator $\Delta_{\phi}|_{W^{2,2}_{0,\phi}(\Omega)}$ in $L^{2}_{\phi}(\Omega)$, which is only discrete and can be deduced from the compactness of the resolvent $\left(-\Delta_{\phi}+I\right)^{-1}$, together with the spectral theory of compact self-adjoint differential operators. See e.g. [38, Subsection 2.2] for details. Combining with the variational principle, one can get two basic facts for the eigenvalue problem (1.11) listed in Section 1.
As explained in [38, Section 3], once the self-adjoint extension of $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$ has been set up as in Theorem 3.1, the heat semigroup $e^{t\Delta_{\phi}}$ related to the Witten-Laplacian $\Delta_{\phi}$ can be well-defined. Then one can get the following properties for the weighted heat kernel by using the tool of heat semigroup.
Lemma 3.2.
(see e.g. [38, Theorem 3.3 in Section 3]) Let $M$ be a complete Riemannian manifold and $\phi$ be a real-valued smooth function on $M$. Denote by $\mathbb{R}^{+}$ the set of all positive real numbers. Then there exists a weighted heat kernel $H^{\phi}(x,y,t)\in C^{\infty}(M\times M\times\mathbb{R}^{+})$ such that
(1) $\lim_{t\rightarrow 0}H^{\phi}(x,y,t)=\delta^{\phi}_{x}(y)$,
(2) $(\Delta_{\phi}-\frac{d}{dt})H^{\phi}=0$,
(3) $H^{\phi}(x,y,t)=H^{\phi}(y,x,t)$,
(4) $H^{\phi}(x,y,t)=\int_{M}H^{\phi}(x,z,t-s)H^{\phi}(z,y,s)d\mu(z)$, with $0<s<t$.
Remark 3.3.
It is easy to see that (1)-(2) in Lemma 3.2 have close relation with the existence of the weight heat kernel. In [38, Theorem 3.3 in Section 3], the property (3) is the symmetry of the weight heat kernel, while the property (4) is the semigroup identity of the weight heat kernel. In [38, Section 3], A. Grigor’yan mentioned at least two approaches to prove the existence of the weight heat kernel — the classical one is the usage of a parametrix of the heat equation, and the other one is the usage of eigenfunction expansion formula in an exhaustion of relatively compact open sets $\{\Omega_{i}\}_{i=1}^{\infty}$ of $M$. For details, one can check e.g. [11, 16, 19, 27, 65] as references.
The following strong maximum (resp., minimum) principle is needed.
Theorem 3.4.
Given a Riemannian manifold $M$ and the Witten-Laplacian $\Delta_{\phi}$, and the associated weighted heat operator $\mathcal{L}^{\phi}=\Delta_{\phi}-\frac{\partial}{\partial t}$, where $\phi$ is a smooth real-valued function on $M$. Let $u(x,t)$ be a bounded continuous function on $M\times[0,T]$, which is $C^{2}$ with respect to the variable $x\in M$, and $C^{1}$ with respect to $t\in[0,T]$, and which satisfies
$\mathcal{L}^{\phi}u\geq 0\qquad(\mathcal{L}^{\phi}u\leq 0)$
on $M\times(0,T)$. If there exists $(x_{0},t_{0})$ in $M\times(0,T]$ such that
$u(x_{0},t_{0})=\sup\limits_{M\times[0,T]}u(x,t)\qquad\left(resp.,~u(x_{0},t_{0})=\inf\limits_{M\times[0,T]}u(x,t)\right),$
then $u|_{M\times[0,t_{0}]}=u(x_{0},t_{0})$.
Proof.
It is easy to check that $\mathcal{L}^{\phi}$ is uniformly parabolic, and then the assertion in Theorem 3.1 directly follows from the standard theory of second-order parabolic partial differential equations (see e.g. [16, p. 180] or [32, Chapter 7]). ∎
We also need:
Lemma 3.5.
For the model space $M^{+}$ (resp., $M^{-}$), if furthermore the potential function $\phi$ has Property 1, its weighted heat kernel $H^{\phi}_{+}(q^{+},z,t)=H^{\phi}_{+}(r_{1},t)$ (resp., $H^{\phi}_{-}(q^{-},z,t)=H^{\phi}_{-}(r_{2},t)$) depends only on variables $r_{1}=d_{M^{+}}(q^{+},z)$ (resp., $r_{2}=d_{M^{-}}(q^{-},z)$) and $t$. Moreover, for all $t>0$, we have
$\frac{\partial}{\partial r_{1}}H^{\phi}_{+}(r_{1},t)<0\qquad\left(\frac{\partial}{\partial r_{2}}H^{\phi}_{-}(r_{2},t)<0\right).$
Proof.
If the potential function $\phi$ has Property 1, then the weighted heat kernel $H_{+}^{\phi}$ (resp, $H^{\phi}_{-}$) on a geodesic ball $\mathscr{B}_{n}(q^{+},\cdot)$ (resp, $\mathscr{B}_{n}(q^{-},\cdot)$) of the model space $M^{+}$ (resp, $M^{-}$) depends on the space variable $r_{1}$ (resp., $r_{2}$) can be observed by the decomposition of the Witten-Laplacian on $\mathscr{B}_{n}(q^{+},\cdot)$ (resp, $\mathscr{B}_{n}(q^{-},\cdot)$) in geodesic spherical coordinates at $q^{+}$ (resp, $q^{-}$) — see the $3^{\rm{rd}}$ line in the proof of Lemma 6.1 for the detailed expression. The monotonicity of $H_{+}^{\phi}$ (resp, $H^{\phi}_{-}$) with respect to the time variable can be obtained by using an almost the same argument to that in the proof of [19, Lemma 2.3]. But there are two differences we wish to point out as follows:
  • In the non-weighted case, [19] has used the fact that for sufficiently small $0<r_{1},t<\epsilon$, the heat kernel $H_{+}(r_{1},t)$ is asymptotical to be
    $H_{+}(r_{1},t)\sim\frac{1}{(4\pi t)^{n/2}}\times\exp\left(\frac{-r_{1}^{2}}{4t}\right)\sum\limits_{i=0}^{\infty}a_{i}(r_{1})t^{i},$
    where $a_{i}(r_{1})$ are smooth functions on $M^{+}$. This directly implies the expression (2.11) in [19]. In the weighted case, [38, Theorem 3.9] tells us that for sufficiently small $0<r_{1},t<\epsilon$, there exists a smooth positive function $u(r_{1})$ such that the weighted heat kernel $H_{+}^{\phi}(r_{1},t)$ satisfies
    $H_{+}^{\phi}(r_{1},t)\sim\frac{1}{(4\pi t)^{n/2}}\times\exp\left(\frac{-r_{1}^{2}}{4t}\right)u(r_{1}),$
    which leads to the truth that the analysis process in the proof of [19, Lemma 2.3] before the equation (2.20) can be transplanted to the non-weighted case without any big obstacle. Same story happens to $H^{\phi}_{-}(r_{2},t)$. By the way, a detailed proof of the above asymptotic expansion for the weighted heat kernel can be found in [10, 36].
  • For the second-order parabolic equation (2.20) in [19] for the quantity $\partial H_{+}/\partial r_{1}$, the coefficient $m(r_{1})$ of its first-order derivative term is given by
    $-(n-1)\frac{f^{\prime}(r_{1})}{f(r_{1})},$
    with $f$ appearing in the metric of the model considered therein. Actually, if one checks the details in [19] carefully, he or she would find that a spherically symmetric manifold must be an open or closed Ricci model with respect to its base point (i.e. our model space used here is more accurate than the ones used in [19]), and then except regularity assumptions, the function $f$ in $m(r_{1})$ of the equation (2.20) in [19] should have the same form as our warping function $f$ here determined by the initial value problem (1.5). Now, as the equation (2.20) in [19], if one attempts to derive an equation for the quantity $\partial H^{\phi}_{+}/\partial r_{1}$, the coefficient of its first-order derivative term would become
    $-\left[(n-1)\frac{f^{\prime}(r_{1})}{f(r_{1})}-\phi^{\prime}(r_{1})\right],$
    which has no effect on the uniform parabolicity of that equation. This implies that for the weighted case, similar to what Cheeger-Yau have done for the equation (2.20) in [19], we can also use the maximum principle for the second-order uniformly parabolic partial differential equations to deal with the equation of $\partial H^{\phi}_{+}(r_{1},t)/\partial r_{1}$, and then the result
    $\frac{\partial}{\partial r_{1}}H^{\phi}_{+}(r_{1},t)<0,\quad\forall t>0$
    follows. Same story happens to $H^{\phi}_{-}(r_{2},t)$.
Our proof is finished. ∎
Now, we have:
Proof of Theorem 1.6.
We wish to use a similar argument to that in the proof of our previous work [49, Theorem 6.6].
Assume that $M^{n}$ has a radial Ricci curvature lower bound $(n-1)\kappa(s)$ w.r.t. some point $q\in M^{n}$, $r_{0}<\min\{\ell(q),l\}$, and the potential function $\phi$ satisfies Property 1. Since the geodesic ball $B(q,r_{0})$ maybe has points on the cut-locus of $q$, which leads to the invalidity of the path of linear transformations $\mathbb{A}$ defined as Section 2, we need to use a limit procedure shown in [19] to avoid this problem. By applying Lemma 3.2, one can get
$H^{\phi}(q,y,t)-H^{\phi}_{-}\left(d_{M^{-}}\left(q^{-},z\right),t\right)=H^{\phi}(q,y,t)-H^{\phi}_{-}\left(d_{M^{n}}\left(q,y\right),t\right)$ (3.2)
$=$ $H^{\phi}(q,y,t)-H^{\phi}_{-}\left(r_{2}\left(q,y\right),t\right)$
$=$ $-\int\limits_{0}^{t}\int\limits_{B(q,r_{0})}\frac{\partial}{\partial s}\left[H^{\phi}_{-}\left(r_{2}\left(q,x\right),t-s\right)\right]H^{\phi}(x,y,s)d\mu(x)ds$
$+\int\limits_{0}^{t}\int\limits_{B(q,r_{0})}H^{\phi}_{-}\left(r_{2}\left(q,x\right),t-s\right)\frac{\partial H^{\phi}}{\partial s}(x,y,s)d\mu(x)ds$
$=$ $-\int\limits_{0}^{t}\int\limits_{B(q,r_{0})}\frac{\partial}{\partial s}\left[H^{\phi}_{-}\left(r_{2}\left(q,x\right),t-s\right)\right]H^{\phi}(x,y,s)e^{-\phi}dv(x)ds$
$+\int\limits_{0}^{t}\int\limits_{B(q,r_{0})}H^{\phi}_{-}\left(r_{2}\left(q,x\right),t-s\right)\frac{\partial H^{\phi}}{\partial s}(x,y,s)e^{-\phi}dv(x)ds.$
For any $\xi\in S^{n-1}_{q}\subseteq T_{q}M^{n}$, let $a(\xi)=\min\{d_{\xi},r_{0}\}$ be defined as in Section 2. It is not hard to see from the definition of $d_{\xi}$ in (2) that $a(\xi)$ is a continuous function on $S^{n-1}_{q}$. As in [18], we can choose a sequence of smooth functions $a_{\epsilon}$ on $S^{n-1}_{q}$, with $a_{\epsilon}(\xi)<a(\xi)$ for any $\xi\in S^{n-1}_{q}$, such that $a_{\epsilon}(\cdot)$ converges uniformly to $a(\cdot)$ as $\epsilon\rightarrow 0$ and moreover the set
$V_{\epsilon}=\{\exp_{q}(t\xi)|t\leq a_{\epsilon}(\xi)\}$
is compact. Clearly, $V_{\epsilon}$ is within the cut locus of $q$. Then it follows that (3.2) becomes
$H^{\phi}(q,y,t)-H^{\phi}_{-}\left(d_{M^{-}}\left(q^{-},z\right),t\right)$ (3.3)
$=$ $H^{\phi}(q,y,t)-H^{\phi}_{-}\left(r_{2}\left(q,y\right),t\right)$
$=$ $\lim\limits_{\epsilon\rightarrow 0}\Bigg\{-\int\limits_{0}^{t}\int\limits_{V_{\epsilon}}\frac{\partial}{\partial s}\left[H^{\phi}_{-}\left(r_{2}\left(q,x\right),t-s\right)\right]H^{\phi}(x,y,s)d\mu(x)ds$
$+\int\limits_{0}^{t}\int\limits_{V_{\epsilon}}H^{\phi}_{-}\left(r_{2}\left(q,x\right),t-s\right)\frac{\partial H^{\phi}}{\partial s}(x,y,s)d\mu(x)ds\Bigg\}$
$=$ $\lim\limits_{\epsilon\rightarrow 0}\Bigg\{-\int\limits_{0}^{t}\int\limits_{V_{\epsilon}}(\Delta_{\phi})_{M^{-}}\left[H^{\phi}_{-}\left(r_{2}\left(q,x\right),t-s\right)\right]H^{\phi}(x,y,s)d\mu(x)ds$
$+\int\limits_{0}^{t}\int\limits_{V_{\epsilon}}H^{\phi}_{-}\left(r_{2}\left(q,x\right),t-s\right)(\Delta_{\phi})_{M^{n}}H^{\phi}(x,y,s)d\mu(x)ds\Bigg\},$
where $(\Delta_{\phi})_{M^{-}}$, $(\Delta_{\phi})_{M^{n}}$ are the Witten-Laplace operators on $M^{-}$ and $M^{n}$, respectively. Applying Green’s formula (3.1), together with either Dirichlet or Neumann boundary condition, it can be deduced from (3.3) that
$H^{\phi}(q,y,t)-H^{\phi}_{-}\left(d_{M^{-}}\left(q^{-},z\right),t\right)$ (3.4)
$=$ $\lim\limits_{\epsilon\rightarrow 0}\Bigg\{\int\limits_{0}^{t}\int\limits_{V_{\epsilon}}\left[(\Delta_{\phi})_{M^{n}}H^{\phi}_{-}\left(r_{2}\left(q,x\right),t-s\right)-(\Delta_{\phi})_{M^{-}}H^{\phi}_{-}\left(r_{2}\left(q,x\right),t-s\right)\right]\times$
$\qquad\qquad\qquad\qquad H^{\phi}(x,y,s)d\mu(x)ds\Bigg\}.$
On the other hand, in the geodesic spherical coordinates near $q$ or $q^{-}$, for function of $r_{2}(q,y)=d_{M^{-}}(q^{-},z)=d_{M^{n}}(q,y)$, one has
$(\Delta_{\phi})_{M^{n}}=\frac{\partial^{2}}{\partial r_{2}^{2}}+\left[\frac{\left(\sqrt{|g|}\right)^{\prime}}{\sqrt{|g|}}-\phi^{\prime}(r_{2})\right]\frac{\partial}{\partial r_{2}}=\frac{\partial^{2}}{\partial r_{2}^{2}}+\left[\frac{\left(J^{n-1}(r_{2},\xi)\right)^{\prime}}{J^{n-1}(r_{2},\xi)}-\phi^{\prime}(r_{2})\right]\frac{\partial}{\partial r_{2}},$
$(\Delta_{\phi})_{M^{-}}=\frac{\partial^{2}}{\partial r_{2}^{2}}+\left[\frac{\left(f^{n-1}(r_{2})\right)^{\prime}}{f^{n-1}(r_{2})}-\phi^{\prime}(r_{2})\right]\frac{\partial}{\partial r_{2}}.$
Substituting the above two equalities into (3.4) and applying Lemma 3.5, we can obtain
$\qquad H^{\phi}(q,y,t)-H^{\phi}_{-}\left(d_{M^{-}}\left(q^{-},z\right),t\right)$ (3.5)
$=\lim\limits_{\epsilon\rightarrow 0}\left\{\int\limits_{0}^{t}\int\limits_{V_{\epsilon}}(n-1)\left[\frac{J^{\prime}(r_{2},\xi)}{J(r_{2},\xi)}-\frac{f^{\prime}(r_{2})}{f(r_{2})}\right]\frac{\partial H^{\phi}_{-}\left(r_{2}\left(q,x\right),t-s\right)}{\partial r_{2}}H^{\phi}(x,y,s)d\mu(x)ds\right\}$
$\geq 0,$
where the fact $d\theta(s,\xi)/ds=(n-1)\left[J(s,\xi)/f(s)\right]^{n-2}\left[J(s,\xi)/f(s)\right]^{\prime}\leq 0$ shown in Theorem 2.6 has also been used. This directly implies $H^{\phi}(q,y,t)\geq H^{\phi}_{-}\left(d_{M^{-}}\left(q^{-},z\right),t\right)$, which is exactly the weighted heat kernel comparison (1.7). When the equality in (1.7) holds at some $(y_{0},t_{0})\in B(q,r_{0})\times(0,\infty)$, by Theorem 3.4 one has $H^{\phi}(q,y,t)=H^{\phi}_{-}\left(d_{M^{-}}\left(q^{-},z\right),t\right)=H^{\phi}(q,y_{0},t_{0})$ on $B(q,r_{0})\times[0,t_{0}]$. Together with (3.5), we know that
$\frac{J^{\prime}(r_{2},\xi)}{J(r_{2},\xi)}-\frac{f^{\prime}(r_{2})}{f(r_{2})}=0$
holds on $B(q,r_{0})$, which implies $J(r_{2},\xi)=f(r_{2})$ holds on $B(q,r_{0})$, and moreover
$\mathrm{vol}(B(q,r_{0}))=\mathrm{vol}(\mathscr{B}_{n}(q^{-},r_{0})).$
Then it follows that $B(q,r_{0})$ is isometric to $\mathscr{B}_{n}(q^{-},r_{0})$ by directly applying Theorem 2.6. This completes the first assertion (1) of Theorem 1.6.
Assume that $M^{n}$ has a radial sectional curvature upper bound $\kappa(s)$ w.r.t. $q\in M^{n}$, $r_{0}<\min\{\mathrm{inj}(q),l\}$, and the potential function $\phi$ satisfies Property 1. Clearly, in this case, the geodesic ball $B(q,r_{0})$ must be within the cut-locus of $q$. By applying Lemma 3.2, Green’s formula (3.1), Lemma 3.5, and using similar calculations to the previous case, we have
$H^{\phi}(q,y,t)-H^{\phi}_{-}\left(d_{M^{+}}\left(q^{+},z\right),t\right)$ (3.6)
$=$ $H^{\phi}(q,y,t)-H^{\phi}_{+}\left(d_{M^{n}}\left(q,y\right),t\right)$
$=$ $H^{\phi}(q,y,t)-H^{\phi}_{+}\left(r_{1}\left(q,y\right),t\right)$
$=$ $-\int\limits_{0}^{t}\int\limits_{B(q,r_{0})}\frac{\partial}{\partial s}\left[H^{\phi}_{+}\left(r_{1}\left(q,x\right),t-s\right)\right]H^{\phi}(x,y,s)d\mu(x)ds$
$+\int\limits_{0}^{t}\int\limits_{B(q,r_{0})}H^{\phi}_{+}\left(r_{1}\left(q,x\right),t-s\right)\frac{\partial H^{\phi}}{\partial s}(x,y,s)d\mu(x)ds$
$=$ $\int\limits_{0}^{t}\int\limits_{B(q,r_{0})}\left[(\Delta_{\phi})_{M^{n}}H^{\phi}_{+}\left(r_{1}\left(q,x\right),t-s\right)-(\Delta_{\phi})_{M^{+}}H^{\phi}_{+}\left(r_{1}\left(q,x\right),t-s\right)\right]\times$
$\qquad\qquad\qquad\qquad H^{\phi}(x,y,s)d\mu(x)ds$
$=$ $\int\limits_{0}^{t}\int\limits_{B(q,r_{0})}\left[\frac{\left(\sqrt{|g|}\right)^{\prime}}{\sqrt{|g|}}-\frac{\left(f^{n-1}(r_{1})\right)^{\prime}}{f^{n-1}(r_{1})}\right]\frac{\partial H^{\phi}_{+}\left(r_{1}\left(q,x\right),t-s\right)}{\partial r_{1}}H^{\phi}(x,y,s)d\mu(x)ds$
$\leq$ $0,$
where the fact
$\left[\frac{\sqrt{|g|}}{f^{n-1}}\right]^{\prime}\geq 0$
shown in Theorem 2.7 has also been used. This implies $H^{\phi}(q,y,t)\leq H^{\phi}_{+}\left(d_{M^{+}}\left(q^{+},z\right),t\right)$, which is exactly the weighted heat kernel comparison (1.8). When the equality in (1.8) holds at some $(y_{0},t_{0})\in B(q,r_{0})\times(0,\infty)$, by Theorem 3.4 one has $H^{\phi}(q,y,t)=H^{\phi}_{+}\left(d_{M^{+}}\left(q^{+},z\right),t\right)=H^{\phi}(q,y_{0},t_{0})$ on $B(q,r_{0})\times[0,t_{0}]$. Together with (3.6), we know that
$\frac{\left(\sqrt{|g|}\right)^{\prime}}{\sqrt{|g|}}=\frac{\left(f^{n-1}(r_{1})\right)^{\prime}}{f^{n-1}(r_{1})}$
holds on $B(q,r_{0})$, which implies $\sqrt{|g|}(r_{1})=f^{n-1}(r_{1})$ holds on $B(q,r_{0})$. By directly applying Theorem 2.7, one knows that $\mathbb{A}(r_{1},\xi)=f(r_{1})I$ holds on $B(q,r_{0})$, which immediately implies that $\mathrm{vol}(B(q,r_{0}))=\mathrm{vol}(\mathscr{B}_{n}(q^{+},r_{0}))$, and $B(q,r_{0})$ is isometric to $\mathscr{B}_{n}(q^{+},r_{0})$. This completes the second assertion (2) of Theorem 1.6. ∎
4 Applications in spectral geometry
Theorem 3.1 tells us that if $\Omega\subseteq M^{n}$ is geodesically complete, with $M^{n}$ a given complete Riemannian $n$-manifold, then $\Delta_{\phi}|_{W^{2,2}_{0,\phi}(\Omega)}$ is a unique self-adjoint extension of $\Delta_{\phi}|_{C^{\infty}_{0}(\Omega)}$, and this fact can be used to prove that the eigenvalue problem (1.11) only has discrete spectrum, which consists of eigenvalues given in the sequence (1.12). Let $\psi_{i}$ be an eigenfunction corresponding to the $i$-th Dirichlet eigenvalue $\lambda_{i,\phi}(\Omega)$ such that $\{\psi_{i}\}_{i=1}^{\infty}$ forms an orthonormal basis in $L^{2}_{\phi}(\Omega)$. From [38, Section 3], one knows that the weighted heat kernel $H^{\phi}(x,y,t)$ on $\Omega$ has the following eigenfunction expansion formula
$H^{\phi}(x,y,t)=\sum\limits_{i=1}^{\infty}e^{-\lambda_{i,\phi}(\Omega)t}\psi_{i}(x)\psi_{i}(y).$ (4.1)
In fact, the formula (4.1) was derived in [38, Section 3] by using several properties of the heat semigroup generated by the Witten-Laplacian $\Delta_{\phi}$. One can also check our recent work [72, Section 5] for a simpler way to get the formula (4.1) and also several other properties of the weighted heat kernel.
Now, by mainly using the formula (4.1), we have:
A proof of Cheng-type eigenvalue comparisons (1.14) and (1.15).
Assume that $M^{n}$ has a radial Ricci curvature lower bound $(n-1)\kappa(s)$ w.r.t. some point $q\in M^{n}$, $r_{0}<\min\{\ell(q),l\}$, and the potential function $\phi$ satisfies Property 1. Then by Theorem 1.6, we have
$H^{\phi}(q,q,t)\geq H^{\phi}_{-}(r_{2}(q,q),t)=H^{\phi}_{-}(0,t)$ (4.2)
for all $t>0$. By applying the expansion formula (4.1), one has
$H^{\phi}(q,q,t)=\sum\limits_{i=1}^{\infty}e^{-\lambda_{i,\phi}t}\psi_{i}^{2}(q),$
$H^{\phi}_{-}(0,t)=\sum\limits_{i=1}^{\infty}e^{-\widetilde{\lambda}_{i,\phi}t}\widetilde{\psi}_{i}^{2}(0),$
where $\lambda_{i,\phi}=\lambda_{i,\phi}(B(q,r_{0}))$, $\widetilde{\lambda}_{i,\phi}=\lambda_{i,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$, and $\psi_{i}$, $\widetilde{\psi}_{i}$ are eigenfunctions corresponding to $\lambda_{i,\phi}(B(q,r_{0}))$, $\lambda_{i,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$ respectively. Putting the above two equalities into (4.2) results in
$e^{-\lambda_{1,\phi}t}\left[\psi_{1}^{2}(q)+e^{-(\lambda_{2,\phi}-\lambda_{1,\phi})t}\psi_{2}^{2}(q)+e^{-(\lambda_{3,\phi}-\lambda_{1,\phi})t}\psi_{3}^{2}(q)+\cdots\right]$
$\qquad\qquad\geq e^{-\widetilde{\lambda}_{1,\phi}t}\left[\widetilde{\psi}_{1}^{2}(0)+e^{-(\widetilde{\lambda}_{2,\phi}-\widetilde{\lambda}_{1,\phi})t}\widetilde{\psi}_{2}^{2}(0)+e^{-(\widetilde{\lambda}_{3,\phi}-\widetilde{\lambda}_{1,\phi})t}\widetilde{\psi}_{3}^{2}(0)+\cdots\right],$
which implies
$\psi_{1}^{2}(q)+e^{-(\lambda_{2,\phi}-\lambda_{1,\phi})t}\psi_{2}^{2}(q)+\cdots\geq e^{(\lambda_{1,\phi}-\widetilde{\lambda}_{1,\phi})t}\left[\widetilde{\psi}_{1}^{2}(0)+e^{-(\widetilde{\lambda}_{2,\phi}-\widetilde{\lambda}_{1,\phi})t}\widetilde{\psi}_{2}^{2}(0)+\cdots\right].$ (4.3)
Since $\psi_{1}^{2}(q)>0$, $\widetilde{\psi}_{1}^{2}(0)$, and $\lambda_{i,\phi}>\lambda_{1,\phi}$, $\widetilde{\lambda}_{i,\phi}>\widetilde{\lambda}_{1,\phi}$ for any $i\geq 2$, letting $t\rightarrow\infty$ in (4.3) leads to
$\lambda_{1,\phi}-\widetilde{\lambda}_{1,\phi}\leq 0,$
which is equivalent to
$\lambda_{1,\phi}(B(q,r_{0}))\leq\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right).$
On the other hand, by applying Theorem 1.6 and using a similar argument as above, one can obtain that for $r_{0}<\min\{\mathrm{inj}(q),l\}$, the inequality
$\lambda_{1,\phi}(B(q,r_{0}))\geq\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{+},r_{0})\right).$
holds when $M^{n}$ has a radial sectional curvature upper bound $\kappa(s)$ w.r.t. $q\in M^{n}$, and the potential function $\phi$ satisfies Property 1. This completes the proof. ∎
Remark 4.1.
Similar to what we have pointed out in [49, Remark 6.9], one cannot get the rigidity characterization for the equality in Cheng-type eigenvalue comparisons (1.14) or (1.15) through the approach used in this section. In fact, if $\lambda_{1,\phi}(B(q,r_{0}))=\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$ here, we can only get $\lim_{t\rightarrow\infty}H^{\phi}(q,q,t)=\lim_{t\rightarrow\infty}H^{\phi}_{-}(0,t)$. We are not sure whether there exists some $t_{0}$ such that $H^{\phi}(q,q,t_{0})=H^{\phi}_{-}(0,t_{0})$ or not, which leads to the fact that we cannot use the rigidity characterization for the weighted heat kernel comparison (1.7) in Theorem 1.6. Same story happens to $\lambda_{1,\phi}(B(q,r_{0}))=\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{+},r_{0})\right)$. By the way, the isometric rigidity characterizations for the eigenvalue comparisons (1.14) and (1.15) in Theorems 1.17-1.18 will be proven in Appendix B using another approach.
Except eigenvalue comparisons, the weighted heat kernel also has some other applications in the field of spectral geometry — such as eigenvalue estimates, spectral isoperimetric inequalities, and so on. We wish to refer readers to the survey [38] for this viewpoint.
5 Appendix A
In this appendix, we derive the explicit expression for the weighted heat kernel in Example 1.7, and the method is standard separation of variables. Assume that the weighted heat kernel is of the form
$H^{\phi}(x_{0},y,t)=H^{\phi}(|y-x_{0}|,t)$ $=$ $\varphi(|y-x_{0}|)\psi(t)\times\exp\left(-\frac{|y-x_{0}|^{2}}{4t}\right)$
$=$ $H^{\phi}(r,t)$
$=$ $\varphi(r)\psi(t)\times\exp\left(-\frac{r^{2}}{4t}\right).$
By a direct calculation, one has
$H^{\phi}_{t}=\varphi e^{-\frac{r^{2}}{4t}}\left(\psi_{t}+\psi\frac{r^{2}}{4t}\right),$
$H^{\phi}_{r}=\psi e^{-\frac{r^{2}}{4t}}\left(\varphi_{r}-\varphi\frac{r}{2t}\right),$
$H^{\phi}_{rr}=\psi e^{-\frac{r^{2}}{4t}}\left[\varphi_{rr}-\varphi_{r}\frac{r}{t}+\varphi\left(\frac{r^{2}}{4t^{2}}-\frac{1}{2t}\right)\right].$
So, the weighted heat equation $H^{\phi}_{t}=\Delta_{\phi}H^{\phi}=H^{\phi}_{rr}-\phi_{r}H^{\phi}_{r}$, with $\phi(r)=kr$, $k=\pm 1$, implies
$\varphi\left(\psi_{t}+\psi\frac{r^{2}}{4t}\right)=\psi\left[\varphi_{rr}-\varphi_{r}\frac{r}{t}+\varphi\left(\frac{r^{2}}{4t^{2}}-\frac{1}{2t}\right)\right]-k\psi\left(\varphi_{r}-\varphi\frac{r}{2t}\right).$
It then follows that
$\frac{\psi_{t}}{\psi}=\frac{\varphi_{rr}-k\varphi_{r}}{\varphi}-\frac{r}{2t}\cdot\frac{2\varphi_{r}-k\varphi}{\varphi}-\frac{1}{2t}.$
Hence, there exist constants $C_{1}$, $C_{2}$ such that
$\frac{\varphi_{rr}-k\varphi_{r}}{\varphi}=C_{1},\qquad\frac{(2\varphi_{r}-k\varphi)r}{\varphi}=C_{2},\qquad\frac{\psi_{t}}{\psi}=C_{1}-\frac{C_{2}+1}{2t}.$
Solving the above ODEs yields
$\varphi(r)=C_{3}\cdot\exp\left(\frac{1}{2}kr\right),$
$\psi(t)=C_{4}\cdot\frac{1}{\sqrt{t}}\exp\left(-\frac{4}{t}\right),$
where $C_{3}$ is independent of $r$, and $C_{4}$ is independent of $t$. Using the initial condition
$\lim\limits_{t\rightarrow 0}H^{\phi}(x_{0},y,t)=\delta^{\phi}_{y}(x_{0})$
at the point $x_{0}$, one can get
$C_{3}C_{4}=\frac{1}{2\sqrt{\pi}}e^{\frac{1}{2}kr},$
which implies that $C_{4}$ should depend on $r$.
Summing up, the weighted heat kernel in Example 1.7 is explicitly given by
$H^{\phi}(r,t)=\frac{e^{\pm r}\cdot e^{-t/4}}{(4\pi t)^{1/2}}\times\exp\left(-\frac{r^{2}}{4t}\right),$
which is equivalent to say
$H^{\phi}(x_{0},y,t)=H^{\phi}(|y-x_{0}|,t)=\frac{e^{\pm|y-x_{0}|}\cdot e^{-t/4}}{(4\pi t)^{1/2}}\times\exp\left(-\frac{|y-x_{0}|^{2}}{4t}\right).$
6 Appendix B
In order to prove Theorem 1.17, we need the following property.
Lemma 6.1.
Assume that the potential function $\phi$ has Property 1. The eigenfunction corresponding to the first Dirichlet eigenvalue of the Witten-Laplacian on the geodesic ball $\mathscr{B}_{n}(q^{-},r_{0})$ (resp., $\mathscr{B}_{n}(q^{+},r_{0})$) may be chosen to be nonnegative and is a radial function $\psi(s)$ satisfying $\psi^{\prime}(s)<0$ for $0<s\leq r_{0}<\min\{\ell(q),l\}$, with $\ell(q)$ defined as in (2.7) (resp., for $0<s\leq r_{0}<\min\{\mathrm{inj}(q),l\}$, with $\mathrm{inj}(q)$ defined by (2.5)).
Proof.
The Witten-Laplacian on $\mathscr{B}_{n}(q^{-},r_{0})$ in geodesic spherical coordinates at $q^{-}$ is given by
$\Delta_{\phi}=\frac{d^{2}}{ds^{2}}+(n-1)\frac{f^{\prime}(s)}{f(s)}\frac{d}{ds}+\frac{1}{f^{2}(s)}\Delta_{\mathbb{S}^{n-1}}-\langle\phi^{\prime}(s),\nabla\cdot\rangle,$
where $\Delta_{\mathbb{S}^{n-1}}$ denotes the Laplacian on $\mathbb{S}^{n-1}$ with respect to the round metric. Using a similar argument to Courant’s nodal domain theorem for the Laplacian shown in [16, Chapter I], one knows that the first Dirichlet eigenvalue $\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$ of the Witten-Laplacian is simple (i.e. the eigenspace of this eigenvalue has dimension $1$) and its eigenfunction $\psi$ would not change sign on $\mathscr{B}_{n}(q^{-},r_{0})$. One can also check the proof of our recent work [23, Theorem 1.2] for details if he or she is not familiar with this content. Without loss of generality, one may choose that $\psi$ is nonnegative on $\mathscr{B}_{n}(q^{-},r_{0})$. More precisely, $\psi$ is strictly positive within $\mathscr{B}_{n}(q^{-},r_{0})$ and vanishes on the boundary $\partial\mathscr{B}_{n}(q^{-},r_{0})$. Since the dimension of the eigenspace corresponding to $\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$ is $1$, it follows that $\psi$ is radial and satisfies
$\frac{d^{2}\psi}{ds^{2}}+\left[(n-1)\frac{f^{\prime}(s)}{f(s)}-\phi^{\prime}(s)\right]\frac{d\psi}{ds}+\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)\psi(s)=0,$
i.e.
$\psi^{\prime\prime}(s)+\left[(n-1)\frac{f^{\prime}(s)}{f(s)}-\phi^{\prime}(s)\right]\psi^{\prime}(s)+\lambda_{1,\phi}\psi(s)=0.$ (6.1)
Set $\alpha(s):=e^{\int_{\epsilon}^{s}\left[(n-1)\frac{f^{\prime}(t)}{f(t)}-\phi^{\prime}(t)\right]dt}$, $\epsilon\rightarrow 0^{+}$, and then it is not hard to know that (6.1) becomes
$(\alpha\psi^{\prime})^{\prime}+\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)\alpha\psi=0.$ (6.2)
By a direct calculation, one has from (6.2) that
$(\alpha\psi^{\prime\prime})^{\prime}+\left[\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)+\left(\frac{\alpha^{\prime}}{\alpha}\right)^{\prime}\right]\alpha\psi^{\prime}=0.$
Integrating both sides of (6.2) on the interval $(0,s)$ with $s<r_{0}$, one has
$\alpha(s)\psi^{\prime}(s)=-\lambda_{1,\phi}\int_{0}^{s}\alpha(t)\psi(t)dt,$
and this implies that $\psi^{\prime}(s)|_{(0,r_{0}]}<0$ whenever $\psi(s)|_{(0,r_{0})}>0$ and $\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)>0$. The situation for the eigenfunction on the geodesic ball $\mathscr{B}_{n}(q^{+},r_{0})$ has no essential difference. Our proof is finished. ∎
Remark 6.2.
As maybe indicated before, the reason that we require $0<s\leq r_{0}<\min\{\ell(q),l\}$ (resp., $0<s\leq r_{0}<\min\{\mathrm{inj}(q),l\}$) is that when the equality holds in the Cheng-type eigenvalue comparison (1.14) (resp., (1.15)), the isometric rigidity between geodesic balls $B(q,r_{0})$ and $\mathscr{B}_{n}(q^{-},r_{0})$ (resp., $\mathscr{B}_{n}(q^{+},r_{0})$) can be assured.
Now, together with Lemma 6.1, we wish to prove Theorem 1.17 by using a similar argument to that in the proof of [33, Theorem 3.6].
Proof of Theorem 1.17.
Let $\psi$ be an eigenfunction of the Dirichlet eigenvalue $\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$, which, by Lemma 6.1, is simple, radial and can be chosen to be nonnegative. Since $\psi\circ r$ vanishes on the boundary $\partial B(q,r_{0})$, where as in Section 2, $r(x)=d_{M^{n}}(q,x)$ measures the radial distance from $q$ to any point $x\in\partial B(q,r_{0})$, it is not hard to know that $\psi\circ r\in W^{1,2}_{0,\phi}(B(q,r_{0}))$ and
$\lambda_{1,\phi}(B(q,r_{0}))\leq\frac{\int\limits_{B(q,r_{0})}\langle d\psi\circ r,d\psi\circ r\rangle e^{-\phi}dv}{\int\limits_{B(q,r_{0})}\langle\psi\circ r,\psi\circ r\rangle e^{-\phi}dv}$
by directly using the Rayleigh characterization (1.13).
Using spherical geodesic coordinates centered at $q$ under the integrals, one can obtain
$\int\limits_{B(q,r_{0})}\langle d\psi\circ r,d\psi\circ r\rangle e^{-\phi}dv=\int\limits_{\xi\in S_{q}^{n-1}}\left[\int\limits_{0}^{a(\xi)}\left(\frac{d\psi}{ds}\right)^{2}f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds\right]d\sigma$
and
$\int\limits_{B(q,r_{0})}\langle\psi\circ r,\psi\circ r\rangle e^{-\phi}dv=\int\limits_{\xi\in S_{q}^{n-1}}\left[\int\limits_{0}^{a(\xi)}\psi^{2}(s)f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds\right]d\sigma,$
where as in Section 2, $d\sigma$ is the canonical measure of $\mathbb{S}^{n-1}\equiv S_{q}^{n-1}$, $a(\xi)=\min\{d_{\xi},r_{0}\}$, and $\theta(s,\xi)=\left[J(s,\xi)/f(s)\right]^{n-1}$ on $M^{n}\setminus Cut(q)$ with $f(s)$, $J(s,\xi)$ determined by (1.5) and (2.3) respectively.
On the other hand, by a direct calculation we have
$\int\limits_{0}^{a(\xi)}\left(\frac{d\psi}{ds}\right)^{2}f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds=\psi(s)e^{-\phi}\left(\frac{d\psi(s)}{ds}\right)f^{n-1}(s)\theta(s,\xi)\Big|_{0}^{a(\xi)}$-
$\qquad\int\limits_{0}^{a(\xi)}\frac{\psi(s)}{f^{n-1}(s)\theta(s,\xi)e^{-\phi}}\frac{d}{ds}\left[f^{n-1}(s)\theta(s,\xi)e^{-\phi}\frac{d\psi(s)}{ds}\right]f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds\qquad\qquad$ (6.3)
and
$\frac{1}{f^{n-1}(s)\theta(s,\xi)e^{-\phi}}\frac{d}{ds}\left[f^{n-1}(s)\theta(s,\xi)e^{-\phi}\frac{d\psi(s)}{ds}\right]$
$\quad=\frac{d^{2}\psi}{ds^{2}}+\left[(n-1)\frac{f^{\prime}(s)}{f(s)}-\phi^{\prime}(s)+\frac{d\theta(s,\xi)}{ds}\cdot\frac{1}{\theta(s,\xi)}\right]\frac{d\psi}{ds}$
$\quad=\frac{d^{2}\psi}{ds^{2}}+\left\{(n-1)\frac{f^{\prime}(s)}{f(s)}-\phi^{\prime}(s)+(n-1)\frac{f(s)}{J(s,\xi)}\left[\frac{J(s,\xi)}{f(s)}\right]^{\prime}\right\}\frac{d\psi}{ds}.$ (6.4)
By Lemma 6.1, $\psi^{\prime}(s)<0$ for $0<s<r_{0}$. Together with (6.1), one has from (6) that
$\frac{1}{f^{n-1}(s)\theta(s,\xi)e^{-\phi}}\frac{d}{ds}\left[f^{n-1}(s)\theta(s,\xi)e^{-\phi}\frac{d\psi(s)}{ds}\right]$
$\qquad\qquad\qquad\qquad\qquad=-\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)\psi(s)$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+(n-1)\frac{f(s)}{J(s,\xi)}\left[\frac{J(s,\xi)}{f(s)}\right]^{\prime}\frac{d\psi}{ds}$
$\qquad\qquad\qquad\qquad\qquad\geq-\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)\psi(s)\qquad$ (6.5)
by using the fact $d\theta(s,\xi)/ds=(n-1)\left[J(s,\xi)/f(s)\right]^{n-2}\left[J(s,\xi)/f(s)\right]^{\prime}\leq 0$ shown in Theorem 2.6. Substituting (6) into (6) yields
$\int\limits_{0}^{a(\xi)}\left(\frac{d\psi}{ds}\right)^{2}f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds=\psi(a(\xi))\psi^{\prime}(a(\xi))f^{n-1}(a(\xi))\theta(a(\xi),\xi)e^{-\phi(a(\xi))}$
$+\int\limits_{0}^{a(\xi)}\psi^{2}(s)\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)\cdot f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds$
$\leq\int\limits_{0}^{a(\xi)}\psi^{2}(s)\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)\cdot f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds,$ (6.6)
where the facts $\psi^{\prime}(0)=0$ and $\psi(a(\xi))\psi^{\prime}(a(\xi))\leq 0$ have been used. Therefore, it follows from (6) that
$\int\limits_{\xi\in S_{q}^{n-1}}\left[\int\limits_{0}^{a(\xi)}\left(\frac{d\psi}{ds}\right)^{2}f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds\right]d\sigma\leq$
$\qquad\qquad\qquad\int\limits_{\xi\in S_{q}^{n-1}}\left[\int\limits_{0}^{a(\xi)}\psi^{2}(s)\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)\cdot f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds\right],$
which directly implies $\lambda_{1,\phi}(B(q,r_{0}))\leq\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$. When equality holds, one has $a(\xi)=r_{0}$ for almost all $\xi\in S^{n-1}_{q}\equiv\mathbb{S}^{n-1}$. Hence, $a(\xi)\equiv r_{0}$ for all $\xi$, and from which we have $J(s,\xi)=f(s)$, and then $\mathrm{vol}(B(q,r_{0}))=\mathrm{vol}(\mathscr{B}_{n}(q^{-},r_{0}))$. This implies from Theorem 2.6 that $B(q,r_{0})$ is isometric to $\mathscr{B}_{n}(q^{-},r_{0})$. ∎
Now, we wish to prove Theorem 1.18. However, we need first the following fact, which can be seen as a weighted version of Barta’s lemma (see [8]).
Lemma 6.3.
Let $\Omega$ be a normal domain in a Riemannian manifold with density $e^{-\phi}$, where $\phi$ is a smooth real-valued function defined on this manifold, and let $h\in C^{2}(\Omega)\cap C^{0}(\bar{\Omega})$, with $h|_{\Omega}>0$ and $h|_{\partial\Omega}=0$. Then
$\inf_{\Omega}\left(\frac{\Delta_{\phi}h}{h}\right)\leq-\lambda_{1,\phi}(\Omega)\leq\sup_{\Omega}\left(\frac{\Delta_{\phi}h}{h}\right),$
where $\lambda_{1,\phi}(\Omega)$ denotes the lowest Dirichlet eigenvalue of the Witten-Laplacian on the domain $\Omega$.
Proof.
Let $\Psi$ be an eigenfunction corresponding to $\lambda_{1,\phi}(\Omega)$, with $\Psi|_{\Omega}>0$ and $\Psi|_{\partial\Omega}=0$. This choice of the eigenfunction $\Psi$ can be assured by [23, Theorem 1.2], i.e. the number of nodal domains of the eigenfunction $\Psi$ is $1$, which implies $\Psi$ would not change sign on $\Omega$. Set
$\mathcal{F}:=\Psi-h,$
and then it follows from the fact $\Delta_{\phi}\Psi+\lambda_{1,\phi}\Psi=0$ that
$-\lambda_{1,\phi}(\Omega)$ $=$ $-\lambda_{1,\phi}=\frac{\Delta_{\phi}\Psi}{\Psi}=\frac{\Delta\Psi}{\Psi}-\frac{\langle\nabla\phi,\nabla\Psi\rangle}{\Psi}$ (6.7)
$=$ $\frac{\Delta(\mathcal{F}+h)}{\mathcal{F}+h}-\frac{\langle\nabla\phi,\nabla(\mathcal{F}+h)\rangle}{\mathcal{F}+h}$
$=$ $\frac{\Delta h}{h}+\frac{h\Delta\mathcal{F}-\mathcal{F}\Delta h}{h(h+\mathcal{F})}-\frac{\langle\nabla\phi,\nabla h\rangle}{h}+\frac{\mathcal{F}\langle\nabla\phi,\nabla h\rangle-h\langle\nabla\phi,\nabla\mathcal{F}\rangle}{h(h+\mathcal{F})}$
$=$ $\frac{\Delta_{\phi}h}{h}+\frac{h\Delta_{\phi}\mathcal{F}-\mathcal{F}\Delta_{\phi}h}{h(\mathcal{F}+h)}.$
Since $h|_{\Omega}>0$, $h|_{\partial\Omega}=0$, $h(\mathcal{F}+h)|_{\Omega}>0$ and $h(\mathcal{F}+h)|_{\partial\Omega}=0$, it is not hard to know
$\int_{\Omega}(h\Delta_{\phi}\mathcal{F}-\mathcal{F}\Delta_{\phi}h)e^{-\phi}dv=0,$
which implies that in the domain $\Omega$, the function $h\Delta_{\phi}\mathcal{F}-\mathcal{F}\Delta_{\phi}h$ would have both positive and negative values at different parts. Then it is easy to know from (6.7) that
$\lambda_{1,\phi}(\Omega)\leq\sup_{\Omega}\left(\frac{\Delta_{\phi}h}{h}\right)$
holds if $h\Delta_{\phi}\mathcal{F}-\mathcal{F}\Delta_{\phi}h$ takes negative values, while
$\inf_{\Omega}\left(\frac{\Delta_{\phi}h}{h}\right)\leq-\lambda_{1,\phi}(\Omega)$
holds if $h\Delta_{\phi}\mathcal{F}-\mathcal{F}\Delta_{\phi}h$ takes positive values. This completes the proof of Lemma 6.3. ∎
Remark 6.4.
From the above proof of Lemma 6.3, one easily knows that its assertion is valid without requiring that the potential function $\phi$ has Property 1.
Proof of Theorem 1.18.
We shall use a similar argument to that in the proof of [33, Theorem 4.4]. Let $\psi(s):[0,r_{0}]\rightarrow[0,\infty)$ be a nonnegative radial eigenfunction of $\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{+},r_{0})\right)$ of $M^{+}$. By Lemma 6.1, one knows that the eigenfunction $\psi$ satisfies (6.1), with $\psi^{\prime}(0)=\psi(r_{0})=0$, $\psi\geq 0$ on $0,r_{0})$, and $\frac{d\psi}{ds}<0$ on $(0,r_{0})$. Define a function $\mathcal{F}:\overline{B(q,r_{0})}\rightarrow[0,\infty)$ by
$\mathcal{F}(\exp_{q}(s\xi))=\psi(s)$
for $(s,\xi)\in[0,r_{0})\times S_{q}^{n-1}$. Then by a straightforward calculation in geodesic spherical coordinates around $q$ and using the characterization (1.13), one can get
$\frac{\Delta_{\phi}\mathcal{F}}{\mathcal{F}}(\exp_{q}(s\xi))\leq\frac{1}{\psi}\left[\psi^{\prime\prime}(s)+(n-1)\frac{f^{\prime}(s)}{f(s)}\psi^{\prime}(s)-\phi^{\prime}(s)\psi^{\prime}(s)\right]=-\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{+},r_{0})\right),$
which, together with Lemma 6.3, implies
$-\lambda_{1,\phi}(B(q,r))\leq\sup\frac{\Delta_{\phi}\mathcal{F}}{\mathcal{F}}(\exp_{q}(s\xi))\leq-\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{+},r_{0})\right),$
i.e. $\lambda_{1,\phi}(B(q,r))\geq\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{+},r_{0})\right)$. When the equality holds, the isometric rigidity between $B(q,r_{0})$ and $\mathscr{B}_{n}(q^{+},r_{0})$ can be obtained by directly using Theorem 2.7. ∎
As inspired by [49], we found that the assertions in Theorem 1.17 can be improved to the case of weighted $p$-Laplacian $\Delta_{p,\phi}$, $1<p<\infty$. In order to get that improvement, we need first the following fact.
Lemma 6.5.
Assume that the potential function $\phi$ has Property 1. The eigenfunction corresponding to the first Dirichlet eigenvalue of the weighted $p$-Laplacian ($1<p<\infty$) on the geodesic ball $\mathscr{B}_{n}(q^{-},r_{0})$ may be chosen to be nonnegative and is a radial function $\psi(s)$ satisfying $\psi^{\prime}(s)<0$ for $0<s\leq r_{0}<\min\{\ell(q),l\}$, with $\ell(q)$ defined as in (2.7). The conclusion is still valid for the eigenfunction belonging to the first Dirichlet eigenvalue of the weighted $p$-Laplacian on the geodesic ball $\mathscr{B}_{n}(q^{+},r_{0})$ for $0<s\leq r_{0}<\min\{\mathrm{inj}(q),l\}$, with $\mathrm{inj}(q)$ defined by (2.5).
Proof.
As mentioned at the end of Section 2, by making suitable adjustments to the arguments in [2, 3, 9, 35, 45] and using the variational principle, some basic properties can be obtained, including:
  • The first Dirichlet eigenvalue $\lambda_{1,p}^{\phi}(\Omega)$ is simple, isolated, and eigenfunctions associated with $\lambda_{1,p}^{\phi}(\Omega)$ do not change sign.
Without loss of generality, one may choose that $\psi$ is nonnegative on $\mathscr{B}_{n}(q^{-},r_{0})$. The first assertion of Lemma 6.5 follows directly by choosing $\Omega=\mathscr{B}_{n}(q^{-},r_{0})$. If furthermore $\phi$ has Property 1, then the weighted $p$-Laplacian $\Delta_{p,\phi}$ on $\mathscr{B}_{n}(q^{-},r_{0})$ in geodesic spherical coordinates at $q^{-}$ is given by
$\Delta_{p,\phi}=|\nabla(\cdot)|^{p-2}\frac{d^{2}}{ds^{2}}+\frac{d}{ds}(|\nabla(\cdot)|^{p-2})\frac{d}{ds}+(n-1)\frac{f^{\prime}(s)}{f(s)}|\nabla(\cdot)|^{p-2}\frac{d}{ds}+\frac{1}{f^{2}(s)}\Delta_{p,\mathbb{S}^{n-1}}$
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad\qquad-|\nabla(\cdot)|^{p-2}\langle\phi^{\prime}(s),\nabla\cdot\rangle,$
where $\Delta_{p,\mathbb{S}^{n-1}}$ denotes the $p$-Laplacian on $\mathbb{S}^{n-1}$ with respect to the round metric. Since $\lambda_{1,p}^{\phi}(\mathscr{B}_{n}(q^{-},r_{0}))$ is simple, isolated, which implies that the dimension of the eigenspace corresponding to $\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$ is $1$, it follows that $\psi$ is radial and satisfies
$(p-1)|\psi^{\prime}(s)|^{p-2}\psi^{\prime\prime}(s)+\left[(n-1)\frac{f^{\prime}(s)}{f(s)}-\phi^{\prime}(s)\right]|\psi^{\prime}(s)|^{p-2}\psi^{\prime}(s)$ (6.8)
$\qquad\qquad\qquad\qquad\qquad\qquad\qquad+\lambda^{\phi}_{1,p}(\mathscr{B}_{n}(q^{-},r_{0}))|\psi(s)|^{p-2}\psi(s)=0$
and the boundary condition $\psi^{\prime}(0)=0$ and $\psi(r_{0})=0$. Set $\alpha(s)=e^{\int_{\epsilon}^{s}\left[(n-1)\frac{f^{\prime}(t)}{f(t)}-\phi^{\prime}(t)\right]dt}$, $\epsilon\rightarrow 0^{+}$, and then it is not hard to know that (6.8) becomes
$(|\psi^{\prime}|^{p-2}\alpha\psi^{\prime})^{\prime}+\lambda^{\phi}_{1,p}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)\alpha\psi|\psi|^{p-2}=0.$ (6.9)
Integrating both sides of (6.9) on the interval $(0,s)$ with $s<r_{0}$, one has
$\alpha(s)|\psi^{\prime}(s)|^{p-2}\psi^{\prime}(s)=-\lambda^{\phi}_{1,p}\int_{0}^{s}\alpha(t)|\psi(t)|^{p-2}\psi(t)dt,$
and this implies that $\psi^{\prime}(s)|_{(0,r_{0}]}<0$ whenever $\psi(s)|_{(0,r_{0})}>0$ and $\lambda^{\phi}_{1,p}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)>0$. ∎
We have the following Cheng-type eigenvalue comparison theorem of the weighted $p$-Laplacian.
Theorem 6.6.
Assume that an $n$-dimensional ($n\geq 2$) complete Riemannian manifold $M^{n}$ has a radial Ricci curvature lower bound $(n-1)\kappa(r)$ w.r.t. $q\in M^{n}$, where $r=d_{M^{n}}(q,\cdot)$ denotes the Riemannian distance from $q$. Assume that the potential function $\phi$ has Property 1. Then for $1<p<\infty$ and $r_{0}<\min\{\ell(q),l\}$ with $\ell(q)$ defined as in (2.7), one has
$\lambda^{\phi}_{1,p}(B(q,r_{0}))\leq\lambda^{\phi}_{1,p}\left(\mathscr{B}_{n}(q^{-},r_{0})\right),$ (6.10)
and the equality holds if and only if $B(q,r_{0})$ is isometric to $\mathscr{B}_{n}(q^{-},r_{0})$.
Proof.
We wish to use a similar argument to that in the proof of [49, Theorem 3.2]. Let $\psi$ be an eigenfunction of the Dirichlet eigenvalue $\lambda^{\phi}_{1,p}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$, which, by Lemma 6.5, is simple, radial and can be chosen to be nonnegative. Since $\psi\circ r$ vanishes on the boundary $\partial B(q,r_{0})$, where $r(x)=d_{M^{n}}(q,x)$ measures the radial distance from $q$ to any point $x\in\partial B(q,r_{0})$, it is not hard to know that $\psi\circ r\in W^{1,p}_{0,\phi}(B(q,r_{0}))$ and
$\lambda^{\phi}_{1,p}(B(q,r_{0}))\leq\frac{\int\limits_{B(q,r_{0})}|d\psi\circ r|^{p}\cdot e^{-\phi}dv}{\int\limits_{B(q,r_{0})}|\psi\circ r|^{p}\cdot e^{-\phi}dv}$
by using the characterization (2.19).
Similar to those calculations in the proof of Theorem 1.17, one can get, by using spherical geodesic coordinates centered at $q$ under the integrals, that
$\int\limits_{B(q,r_{0})}|d\psi\circ r|^{p}\cdot e^{-\phi}dv=\int\limits_{\xi\in S_{q}^{n-1}}\left[\int\limits_{0}^{a(\xi)}|\psi^{\prime}(s)|^{p}f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds\right]d\sigma$
and
$\int\limits_{B(q,r_{0})}|\psi\circ r|^{p}\cdot e^{-\phi}dv=\int\limits_{\xi\in S_{q}^{n-1}}\left[\int\limits_{0}^{a(\xi)}|\psi(s)|^{p}f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds\right]d\sigma.$
On the other hand, since $f(s)>0$ on $(0,r_{0})$ and by Lemma 6.5, $\psi(s)|_{(0,r_{0})}\geq 0$ and $\psi^{\prime}(s)|_{(0,r_{0}]}<0$, it follows by a direct calculation that
$\int\limits_{0}^{a(\xi)}|\psi^{\prime}(s)|^{p}f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds=-\psi(s)|\psi^{\prime}(s)|^{p-1}e^{-\phi}f^{n-1}(s)\theta(s,\xi)\Big|_{0}^{a(\xi)}+$ (6.11)
$\int\limits_{0}^{a(\xi)}\frac{\psi(s)}{f^{n-1}(s)\theta(s,\xi)e^{-\phi}}\frac{d}{ds}\left[|\psi^{\prime}(s)|^{p-1}f^{n-1}(s)\theta(s,\xi)e^{-\phi}\frac{d\psi(s)}{ds}\right]f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds\qquad\quad$
and
$\frac{1}{f^{n-1}(s)\theta(s,\xi)e^{-\phi}}\frac{d}{ds}\left[|\psi^{\prime}(s)|^{p-1}f^{n-1}(s)\theta(s,\xi)e^{-\phi}\right]$
$\quad=-|\psi^{\prime}(s)|^{p-2}\left\{(p-1)\psi^{\prime\prime}(s)+\left[(n-1)\frac{f^{\prime}(s)}{f(s)}-\phi^{\prime}(s)+\frac{d\theta(s,\xi)}{ds}\cdot\frac{1}{\theta(s,\xi)}\right]\psi^{\prime}(s)\right\}$
$\quad=-|\psi^{\prime}(s)|^{p-2}\bigg\{(p-1)\psi^{\prime\prime}(s)+\left[(n-1)\frac{f^{\prime}(s)}{f(s)}-\phi^{\prime}(s)+(n-1)\frac{f(s)}{J(s,\xi)}\left(\frac{J(s,\xi)}{f(s)}\right)^{\prime}\right]$
$\qquad\qquad\qquad\qquad\qquad\cdot\psi^{\prime}(s)\bigg\}.$ (6.12)
Together with (6.8), one has from (6) that
$\frac{1}{f^{n-1}(s)\theta(s,\xi)e^{-\phi}}\frac{d}{ds}\left[|\psi^{\prime}(s)|^{p-1}f^{n-1}(s)\theta(s,\xi)e^{-\phi}\right]$ $=$ $\lambda^{\phi}_{1,p}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)|\psi(s)|^{p-2}\psi(s)$
$-|\psi^{\prime}(s)|^{p-2}(n-1)\frac{f(s)}{J(s,\xi)}\left[\frac{J(s,\xi)}{f(s)}\right]^{\prime}\psi^{\prime}(s)$
$\leq$ $\lambda_{1,p}^{\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)|\psi(s)|^{p-2}\psi(s)\qquad$
by using the fact $d\theta(s,\xi)/ds=(n-1)\left[J(s,\xi)/f(s)\right]^{n-2}\left[J(s,\xi)/f(s)\right]^{\prime}\leq 0$ shown in Theorem 2.6. Substituting the above inequality into (6.11) results in
$\int\limits_{0}^{a(\xi)}|\psi^{\prime}(s)|^{p}f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds=-\psi(a(\xi))|\psi^{\prime}(a(\xi))|^{p-1}f^{n-1}(a(\xi))\theta(a(\xi),\xi)e^{-\phi(a(\xi))}$
$+\int\limits_{0}^{a(\xi)}|\psi(s)|^{p}\lambda^{\phi}_{1,p}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)\cdot f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds$
$\leq\int\limits_{0}^{a(\xi)}|\psi(s)|^{p}\lambda^{\phi}_{1,p}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)\cdot f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds,\qquad$ (6.13)
where the facts $\psi^{\prime}(0)=0$ and $\psi(a(\xi))\geq 0$ have been used. Therefore, it follows from (6) that
$\int\limits_{\xi\in S_{q}^{n-1}}\left[\int\limits_{0}^{a(\xi)}|\psi^{\prime}(s)|^{p}f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds\right]d\sigma\leq$
$\qquad\qquad\qquad\int\limits_{\xi\in S_{q}^{n-1}}\left[\int\limits_{0}^{a(\xi)}|\psi(s)|^{p}\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)\cdot f^{n-1}(s)\theta(s,\xi)e^{-\phi}ds\right],$
which directly implies $\lambda^{\phi}_{1,p}(B(q,r_{0}))\leq\lambda^{\phi}_{1,p}\left(\mathscr{B}_{n}(q^{-},r_{0})\right)$. When equality holds, one has $a(\xi)=r_{0}$ for almost all $\xi\in S^{n-1}_{q}\equiv\mathbb{S}^{n-1}$. Hence, $a(\xi)\equiv r_{0}$ for all $\xi$, and from which we have $J(s,\xi)=f(s)$, and then $\mathrm{vol}(B(q,r_{0}))=\mathrm{vol}(\mathscr{B}_{n}(q^{-},r_{0}))$. This implies from Theorem 2.6 that $B(q,r_{0})$ is isometric to $\mathscr{B}_{n}(q^{-},r_{0})$. ∎
Remark 6.7.
It is easy to see that if $p=2$, then the weighted $p$-Laplacian $\Delta_{p,\phi}$ would degenerate into the Witten-Laplacian $\Delta_{\phi}$ directly, and correspondingly, Theorem 6.6 becomes Theorem 1.17 exactly. However, if one checks the end of Section 2 carefully, he or she would find that the spectral structure of the eigenvalue problem (2.18) is much complicated than the one of the eigenvalue problem (1.11) — the operator $\Delta_{p,\phi}$ in (2.18) has a discrete spectrum and it might also has an essential spectrum, which deeply depends on the geometry and topology of the domain $\Omega$, but the operator $\Delta_{\phi}$ in (1.11) definitely only has a discrete spectrum and each eigenvalue in this discrete spectrum can be characterized clearly.
Remark 6.8.
Denote by $\lambda_{k,p}(\cdot)$ the $k$-th Dirichlet eigenvalue of the $p$-Laplacian $\Delta_{p}$ on a given geometric object. Clearly, if $\phi=const.$, then (6.10) in Theorem 6.6 would reduce to
$\lambda_{1,p}(B(q,r_{0}))\leq\lambda_{1,p}\left(\mathscr{B}_{n}(q^{-},r_{0})\right),$ (6.14)
and the equality holds if and only if $B(q,r_{0})$ is isometric to $\mathscr{B}_{n}(q^{-},r_{0})$. This is exactly the assertions in [49, Theorem 3.2]. If furthermore the curvature assumption was strengthened to be $\mathrm{Ric}(M^{n})\geq(n-1)K$ for some constant $K$, then the eigenvalue comparison (6.14) would reduce to
$\lambda_{1,p}(B(q,r_{0}))\leq\lambda_{1,p}\left(\mathcal{B}_{n}(K,r_{0})\right),$ (6.15)
and the equality holds if and only if $B(q,r_{0})$ is isometric to $\mathcal{B}_{n}(K,r_{0})$. This eigenvalue comparison (7.16) and related rigidity result have already been obtained separately by A. M. Matei [56] and H. Takeuchi [66].
As we know, for an arbitrary point $q\in\mathbb{S}^{n-1}$, its cut locus is a single point set and exactly consists of the antipodal point $\widehat{q}$ of $q$. For a sufficiently small positive constant $\epsilon$, the geodesic ball $B(q,\pi-\epsilon)$ tends to $\mathbb{S}^{n-1}\setminus\{\widehat{q}\}$ as $\epsilon\rightarrow 0^{+}$, which leads to a fact that as $\epsilon\rightarrow 0^{+}$, the first Dirichlet eigenvalue $\lambda_{1}(B(q,\pi-\epsilon))$ would tend to the first closed eigenvalue of the Laplacian on $\mathbb{S}^{n-1}$, that is,
$\lim\limits_{\epsilon\rightarrow 0^{+}}\lambda_{1}(B(q,\pi-\epsilon))=0.$
This spectral asymptotical property on the unit Euclidean (n1)(n-1)-sphere $\mathbb{S}^{n-1}$ has been extended to the first Dirichlet eigenvalue of the Laplacian on “closed” spherically symmetric manifolds (see [33, lemma2.5]), and later to the first Dirichlet eigenvalue of the $p$-Laplacian ($1<p<\infty$) on “closed” spherically symmetric manifolds (see [49, Lemma 2.2]). It is natural to ask whether or not these extensions can be further improved to the case of weighted $p$-Laplacian. In fact, by using the same cut-off continuous functions as those two in the proof of [49, Lemma 2.2], one has:
Lemma 6.9.
Assume that $M^{\ast}$ is a generalized space form $0,l)\times_{f}\mathbb{S}^{n-1}$ (with $q^{\ast}$ as its base point) and $\phi$ is a smooth real-valued function defined on $M^{\ast}$, where $f\in C^{2}([0,l))$ and $C^{3}$ at $s=0$, $f(0)=f^{\prime\prime}(0)=0$, $f^{\prime}(0)=1$, closing at $s=l$, i.e. $f(l)=0$. We have
(I) in case $n=2$, if for some $\varepsilon>0$, $f\in C^{1}([0,l+\varepsilon))$, then $\lim_{r\rightarrow l^{-}}\lambda_{1,p}^{\phi}\left(\mathscr{B}_{n}(q^{\ast},r)\right)=0$ with $1<p\leq 2$;
(II) in case $n\geq 3$, if for some $\varepsilon>0$, $f\in C^{2}([0,l+\varepsilon))$, then $\lim_{r\rightarrow l^{-}}\lambda_{1,p}^{\phi}\left(\mathscr{B}_{n}(q^{\ast},r)\right)=0$ with $1<p<3$.
The reason why the same cut-off continuous functions still work in the weighted case is that $e^{-\phi}$ is always bounded on the geodesic ball $\mathscr{B}_{n}(q^{\ast},r)\subseteq M^{\ast}$, which leads to the result that all integrals appeared in the proof of [49, Lemma 2.2] (i.e. corresponding to the non-weighted case) have no essential difference with those would appear in the weighted case.
It is interesting that nearly all the spectral results in this paper need radial Property 1 for the potential function, but this assumption is not required to get the spectral asymptotical result in Lemma 6.9. This invokes us to consider the following question:
Open problem. What kind of assumptions should be imposed on the potential function $\phi$ for transplanting classical spectral results of the Laplacian or the $p$-Laplacian to the weighted case? What is the weakest assumption in a transplanting process?
7 Appendix C
For a given Riemannian $n$-manifold $M^{n}$ and a point $q\in M^{n}$, similar to the argument for the $p$-Laplacian in our previous work [29, Section 1] or [55, Section 1], it is not hard to get that the limit $\lambda_{1,p}^{\phi}(M^{n}):=\lim\limits_{r_{0}\rightarrow\infty}B(q,r_{0})$ exists and is independent of the choice of $q$. Based on this fact, it is also not hard to know that the fundamental tone of the weighted $p$-Laplacian $\Delta_{p,\phi}$ on $\Omega$ can be well-defined as follows
$\left(\lambda_{1,p}^{\phi}(\Omega)\right)^{\ast}:=\inf\left\{\frac{\int_{\Omega}|u|^{p}e^{-\phi}dv}{\int_{\Omega}|\nabla u|^{p}e^{-\phi}dv}=\frac{\int_{\Omega}|u|^{p}d\mu}{\int_{\Omega}|\nabla u|^{p}d\mu}\Big|u\neq 0,~u\in W^{1,p}_{0,\phi}(\Omega)\right\},$ (7.1)
with notations having the same meanings as before. If $\Omega$ is unbounded, then $(\lambda_{1,p}^{\phi}(\Omega))^{\ast}$ coincides with the infimum of the spectrum $\Sigma\subseteq[0,\infty)$ of the unique self-adjoint extension of the weighted $p$-Laplacian $\Delta_{p,\phi}$ acting on $C^{\infty}_{0}(\Omega)$, which is also denoted by $\Delta_{p,\phi}$. If $\Omega$ has compact closure and boundary $\partial\Omega$, from the characterization (2.19) one knows that $(\lambda_{1,p}^{\phi}(\Omega))^{\ast}$ equals the first Dirichlet eigenvalue $\lambda_{1,p}^{\phi}(\Omega)$ of the weighted $p$-Laplacian, i.e. $(\lambda_{1,p}^{\phi}(\Omega))^{\ast}=\lambda_{1,p}^{\phi}(\Omega)$.
First, we need the following fact, which can be seen as a nonlinear attempt of Lemma 6.3.
Lemma 7.1.
Let $\Omega$ be a domain in a Riemannian manifold with density $e^{-\phi}$, where $\phi$ is a smooth real-valued function defined on this manifold. Assume that $h\in C^{1+\alpha}(\Omega)\cap C^{0}(\bar{\Omega})$ with $\Delta_{p,\phi}h\in C^{0}(\Omega)$ and $h|_{\Omega}>0$, where $\alpha\in(0,1)$. Then the fundamental tone $(\lambda_{1,p}^{\phi}(\Omega))^{\ast}$ of the weighted $p$-Laplacian ($1<p<\infty$) satisfies
$\left(\lambda_{1,p}^{\phi}(\Omega)\right)^{\ast}\geq\inf\limits_{\Omega}\left\{-\frac{\Delta_{p,\phi}h}{h^{p-1}}\right\}.$ (7.2)
Moreover, if $\Omega$ is bounded, then $(\lambda_{1,p}^{\phi}(\Omega))^{\ast}$ becomes the lowest Dirichlet eigenvalue $\lambda_{1,p}^{\phi}(\Omega)$ of the weighted $p$-Laplacian on $\Omega$, and the equality in (7.2) holds if and only if $h=\beta u$ for some constant $\beta>0$, where $u\in W^{1,p}_{0,\phi}(\Omega)$ is a nonnegative minimizer of the Rayleigh quotient.
Proof.
Picone’s classical inequality [62, p. 18] states:
  • If $u:\Omega\rightarrow\mathbb{R}$ and $\nu:\Omega\rightarrow\mathbb{R}$ are two differentiable functions such that $u\geq 0$ and $\nu>0$, then in each connected component of $\Omega$ one has
    $|\nabla u|^{2}+\frac{u^{2}}{\nu^{2}}|\nabla\nu|^{2}-\frac{2u}{\nu}\langle\nabla u,\nabla\nu\rangle=|\nabla u|^{2}-\left\langle\nabla\nu,\nabla\left(\frac{u^{2}}{\nu}\right)\right\rangle\geq 0.$ (7.3)
    Moreover, the equality holds if and only if $u$ is proportional to $\nu$.
It is not hard to see that the lower bound estimate
$\inf_{\Omega}\left(\frac{\Delta_{\phi}h}{h}\right)\leq-\lambda_{1,\phi}(\Omega)$
in Lemma 6.3 can be obtained from (7.3) by taking $u$ as a nonnegative eigenfunction of the first Dirichlet eigenvalue $\lambda_{1,\phi}(\Omega)$ of the Witten-Laplacian, $\nu=h$, and then integrating by parts the right hand side of the equation (7.3) over $\Omega$. This observation was inspired by the argument in the proof of [64, Theorem 1.2] with respect to the weighted volume density $d\mu=e^{-\phi}dv$. It is natural to try to improve this observation to the nonlinear case.
In [1], Picone’s result (7.3) was extended to the following:
$|\nabla u|^{p}+(p-1)\frac{u^{p}}{\nu^{p}}|\nabla\nu|^{p}-p\frac{u^{p-1}}{\nu^{p-1}}|\nabla\nu|^{p-2}\langle\nabla u,\nabla\nu\rangle=$
$\qquad\qquad\qquad\qquad\qquad\qquad|\nabla u|^{p}-|\nabla\nu|^{p-2}\left\langle\nabla\nu,\nabla\left(\frac{u^{p}}{\nu^{p-1}}\right)\right\rangle\geq 0,$ (7.4)
with $1<p<\infty$, where the equality in (7) holds if and only if $u$ is proportional to $\nu$ in each connected component of $\Omega$. It is easy to see that if $p=2$ in (7), then it would degenerate into (7.3) exactly.
Let $\Psi\in C^{\infty}_{0}(\Omega)$ with $\Psi\geq 0$, and choose $h\in C^{1+\alpha}(\Omega)$ with $\Delta_{p,\phi}h\in C^{0}(\Omega)$ such that $h>0$ in $\Omega$. Since $\Psi$ has a compact support $\mathrm{supp}(\Psi)$ in $\Omega$ and $h$ is continuous and strictly positive in $\Omega$, there exists some constant $c>0$ such that $\Psi\geq c$ on $\mathrm{supp}(\Psi)\subset\Omega$, and $\Psi^{p}/h^{p-1}$ is well-defined and is contained in $W^{1,p}_{0,\phi}(\Omega)$. So, $\Psi^{p}/h^{p-1}$ can be used as a trial function in (7.1). By choosing $u=\Psi$, $\nu=h$ in (7), and integrating over $\Omega$ with respect to the weighted volume density $d\mu=e^{-\phi}dv$, we can obtain
$0$ $\leq$ $\int_{\Omega}\left(|\nabla\Psi|^{p}+(p-1)\frac{\Psi^{p}}{h^{p}}|\nabla h|^{p}-p\frac{\Psi^{p-1}}{h^{p-1}}|\nabla h|^{p-2}\langle\nabla\Psi,\nabla h\rangle\right)d\mu$ (7.5)
$=$ $\int_{\Omega}\left(|\nabla\Psi|^{p}+(p-1)\frac{\Psi^{p}}{h^{p}}|\nabla h|^{p}-p\frac{\Psi^{p-1}}{h^{p-1}}|\nabla h|^{p-2}\langle\nabla\Psi,\nabla h\rangle\right)e^{-\phi}dv$
$=$ $\int_{\Omega}\left[|\nabla\Psi|^{p}-|\nabla h|^{p-2}\left\langle\nabla h,\nabla\left(\frac{\Psi^{p}}{h^{p-1}}\right)\right\rangle\right]e^{-\phi}dv$
$=$ $\int_{\Omega}|\nabla\Psi|^{p}e^{-\phi}dv-\int_{\Omega}\left\langle|\nabla h|^{p-2}\nabla h,\nabla\left(\frac{\Psi^{p}}{h^{p-1}}\right)\right\rangle e^{-\phi}dv$
$=$ $\int_{\Omega}|\nabla\Psi|^{p}e^{-\phi}dv+\int_{\Omega}\mathrm{div}\left(e^{-\phi}|\nabla h|^{p-2}\nabla h\right)\cdot\frac{\Psi^{p}}{h^{p-1}}dv$
$=$ $\int_{\Omega}|\nabla\Psi|^{p}e^{-\phi}dv+\int_{\Omega}(\Delta_{p,\phi}h)\cdot\frac{\Psi^{p}}{h^{p-1}}e^{-\phi}dv$
$\leq$ $\int_{\Omega}|\nabla\Psi|^{p}e^{-\phi}dv-\inf\limits_{\Omega}\left\{-\frac{\Delta_{p,\phi}h}{h^{p-1}}\right\}\int_{\Omega}\Psi^{p}e^{-\phi}dv,$
which implies
$\inf\limits_{\Omega}\left\{-\frac{\Delta_{p,\phi}h}{h^{p-1}}\right\}\leq\frac{\int_{\Omega}|\nabla\Psi|^{p}e^{-\phi}dv}{\int_{\Omega}\Psi^{p}e^{-\phi}dv}=\frac{\int_{\Omega}|\nabla\Psi|^{p}d\mu}{\int_{\Omega}\Psi^{p}d\mu}$ (7.6)
with $\Psi\in C^{\infty}_{0}(\Omega)$ and $\Psi\geq 0$. Taking the infimum for the Rayleigh quotient
$\frac{\int_{\Omega}|\nabla\Psi|^{p}e^{-\phi}dv}{\int_{\Omega}\Psi^{p}e^{-\phi}dv}$
over $\Psi\in C^{\infty}_{0}(\Omega)\setminus\{0\}$ and $\Psi\geq 0$, together with the fact that $C^{\infty}_{0}(\Omega)$ is dense in the Sobolev space $W^{1,p}_{0,\phi}$, we can get from (7.1) and (7.6) that the estimate (7.2) in Lemma 7.1 holds.
Assume that $\Omega$ is bounded, then one has first that $(\lambda_{1,p}^{\phi}(\Omega))^{\ast}=\lambda_{1,p}^{\phi}(\Omega)$. Correspondingly, the estimate (7.2) becomes
$\lambda_{1,p}^{\phi}(\Omega)\geq\inf\limits_{\Omega}\left\{-\frac{\Delta_{p,\phi}h}{h^{p-1}}\right\}.$ (7.7)
Now, we wish to give a rigidity characterization for the equality in (7.7). To get this purpose, we shall use a similar argument to that in the proof of [64, Theorem 1.2].
We first show the validity of the sufficient condition. If $h=\beta u$ for some constant $\beta>0$, where $u\in W^{1,p}_{0,\phi}(\Omega)$ is a nonnegative minimizer of the Rayleigh quotient, that is,
$\lambda_{1,p}^{\phi}(\Omega)=\frac{\int_{\Omega}|\nabla u|^{p}e^{-\phi}dv}{\int_{\Omega}|u|^{p}e^{-\phi}dv}$
from the characterization (2.19), then one has
$\Delta_{p,\phi}(\beta u)=e^{\phi}\mathrm{div}(e^{-\phi}|\nabla(\beta u)|^{p-2}\nabla(\beta u))=\beta^{p-1}e^{\phi}\mathrm{div}(e^{-\phi}|\nabla u|^{p-2}\nabla u)=\beta^{p-1}\Delta_{p,\phi}u,$
and then in the sense of weak derivatives,
$-\Delta_{p,\phi}h=-\Delta_{p,\phi}(\beta u)=-\beta^{p-1}\Delta_{p,\phi}u=\beta^{p-1}\lambda_{1,p}^{\phi}(\Omega)u^{p-1}=\lambda_{1,p}^{\phi}(\Omega)h^{p-1}$
holds in $\Omega$. This implies
$\lambda_{1,p}^{\phi}(\Omega)=-\frac{\Delta_{p,\phi}h}{h^{p-1}}=\inf\limits_{\Omega}\left\{-\frac{\Delta_{p,\phi}h}{h^{p-1}}\right\}$
with $h=\beta u$ for some constant $\beta>0$.
At the end, we show the validity of the necessary condition. Since $\Omega$ is bounded, the infimum
$\inf\left\{\frac{\int_{\Omega}|u|^{p}e^{-\phi}dv}{\int_{\Omega}|\nabla u|^{p}e^{-\phi}dv}\Big|u\neq 0,~u\in W^{1,p}_{0,\phi}(\Omega)\right\}$
can be achieved by some nonnegative function $u\in W^{1,p}_{0,\phi}(\Omega)$ which is a weak solution to the eigenvalue problem (2.18). In fact, $u$ is an eigenfunction of the first Dirichlet eigenvalue $\lambda_{1,p}^{\phi}(\Omega)$, and it does not change sign in $\Omega$ (see the end of Section 2). Let $\{\Psi_{n}\}_{n=1}^{\infty}$ be a sequence in the space $C^{\infty}_{0}(\Omega)$ such that $\Psi_{n}\geq 0$ and $\Psi_{n}\rightarrow u$ under the norm $\|\cdot\|^{\phi}_{1,p}$ as $n\rightarrow\infty$. Choosing $u=\Psi_{n}$, $\nu=h$ in (7), and using a similar calculating process as that in (7.5), together with the fact $\Psi_{n}^{p}/h^{p-1}\in W^{1,p}_{0,\phi}(\Omega)$, one has
$\int_{\Omega}|\nabla\Psi_{n}|^{p}e^{-\phi}dv-\int_{\Omega}\left(-\frac{\Delta_{p,\phi}h}{h^{p-1}}\right)\Psi_{n}^{p}\cdot e^{-\phi}dv\geq 0.$
Letting $n\rightarrow\infty$ in the above inequality yields
$0\leq\int_{\Omega}|\nabla u|^{p}e^{-\phi}dv-\int_{\Omega}\left(-\frac{\Delta_{p,\phi}h}{h^{p-1}}\right)u^{p}e^{-\phi}dv=\int_{\Omega}|\nabla u|^{p}d\mu-\int_{\Omega}\left(-\frac{\Delta_{p,\phi}h}{h^{p-1}}\right)|u|^{p}d\mu.\qquad$ (7.8)
It follows from (7.8) that
$\inf\limits_{\Omega}\left\{-\frac{\Delta_{p,\phi}h}{h^{p-1}}\right\}\leq\frac{\int_{\Omega}|\nabla u|^{p}d\mu}{\int_{\Omega}|u|^{p}d\mu}.$ (7.9)
Since $u\in W^{1,p}_{0,\phi}(\Omega)$ is a weak solution to the eigenvalue problem (2.18), we can obtain from (2.19) that
$\frac{\int_{\Omega}|u|^{p}e^{-\phi}dv}{\int_{\Omega}|\nabla u|^{p}e^{-\phi}dv}=\frac{\int_{\Omega}|\nabla u|^{p}d\mu}{\int_{\Omega}|u|^{p}d\mu}=\lambda_{1,p}^{\phi}(\Omega).$
Assume that the equality in (7.2) holds, i.e. the necessary condition holds, which means
$\lambda_{1,p}^{\phi}(\Omega)=\inf\limits_{\Omega}\left\{-\frac{\Delta_{p,\phi}h}{h^{p-1}}\right\}.$
Then one gets that the equalities in (7.8) and (7.9) hold simultaneously. Therefore, it follows from the fact
$\int_{\Omega}\left(-\frac{\Delta_{p,\phi}h}{h^{p-1}}\right)|u|^{p}e^{-\phi}dv=\int_{\Omega}|\nabla h|^{p-2}\left\langle\nabla h,\nabla\left(\frac{u^{p}}{h^{p-1}}\right)\right\rangle e^{-\phi}dv$
that the nonnegative integrand
$|\nabla u|^{p}-|\nabla h|^{p-2}\left\langle\nabla h,\nabla\left(\frac{u^{p}}{h^{p-1}}\right)\right\rangle$
should vanish a.e. in $\Omega$, that is to say,
$|\nabla u|^{p}+(p-1)\frac{u^{p}}{h^{p}}|\nabla h|^{p}-p\frac{u^{p-1}}{h^{p-1}}|\nabla h|^{p-2}\langle\nabla u,\nabla h\rangle=0$ (7.10)
holds a.e. in $\Omega$. Set
$\mathcal{X}(x):=\left[|\nabla u|^{p}-p\frac{u^{p-1}}{h^{p-1}}|\nabla u||\nabla h|^{p-2}|\nabla h|+(p-1)\frac{u^{p}}{h^{p}}|\nabla h|^{p}\right](x),\qquad x\in\Omega$
and
$\mathcal{Y}(x):=\left(p\frac{u^{p-1}}{h^{p-1}}|\nabla u||\nabla h|^{p-2}|\nabla h|-p\frac{u^{p-1}}{h^{p-1}}|\nabla h|^{p-2}\langle\nabla u,\nabla h\rangle\right)(x),\qquad x\in\Omega.$
Therefore, the equation (7.10) can be rewritten as
$\mathcal{X}(x)+\mathcal{Y}(x)=0\quad\mathrm{a.e.~in}~\Omega.$ (7.11)
Using Young’s inequality, one has
$|\nabla u|^{p}+(p-1)\frac{u^{p}}{h^{p}}|\nabla h|^{p}$ $=$ $p\left[\frac{|\nabla u|^{p}}{p}+\frac{\left((u|\nabla h|/h)^{p-1}\right)^{p/(p-1)}}{\frac{p}{p-1}}\right]$
$\geq$ $p|\nabla u|\left(\frac{u|\nabla h|}{h}\right)^{p-1},$
which implies $\mathcal{X}\geq 0$ in $\Omega$. By a direct calculation, one can get
$p\frac{u^{p-1}}{h^{p-1}}|\nabla u||\nabla h|^{p-2}|\nabla h|-p\frac{u^{p-1}}{h^{p-1}}|\nabla h|^{p-2}\langle\nabla u,\nabla h\rangle$ $=$ $p\frac{u^{p-1}}{h^{p-1}}|\nabla h|^{p-2}\left(|\nabla u||\nabla h|-\langle\nabla u,\nabla h\rangle\right)$
$\geq$ $0,$
where the last inequality holds because of the Cauchy-Schwarz inequality. So, $\mathcal{Y}\geq 0$ in $\Omega$. Together with (7.11), one easily knows that
$\mathcal{X}(x)=\mathcal{Y}(x)=0\quad\mathrm{a.e.~in}~\Omega.$
From $\mathcal{Y}=0$, one has
$|\nabla u||\nabla h|=\langle\nabla u,\nabla h\rangle,$
which implies that vector fields $\nabla u$ and $\nabla h$ have the same direction a.e. in $\Omega$. That is to say, there exists a nonnegative measurable function $\chi:\Omega\rightarrow[0,\infty)$ such that
$\nabla h=\chi\cdot\nabla u$ (7.12)
a.e. in $\Omega$. Set $\mathcal{D}:=\{x\in\Omega||\nabla u||\nabla h|>0\}$. Since $\mathcal{X}(x)=0$ a.e. in $\Omega$, for a.e. $x\in\mathcal{D}$ one has
$\left(\frac{h}{u|\nabla h|}\right)^{p}\cdot\left[|\nabla u|^{p}-p\frac{u^{p-1}}{h^{p-1}}|\nabla u||\nabla h|^{p-2}|\nabla h|+(p-1)\frac{u^{p}}{h^{p}}|\nabla h|^{p}\right]=0,$
that is,
$\left(\frac{h}{u}\frac{|\nabla u|}{|\nabla h|}\right)^{p}-p\frac{h}{u}\frac{|\nabla u|}{|\nabla h|}+(p-1)=0\quad\mathrm{a.e.~in}~\mathcal{D}.$ (7.13)
Set
$\mathcal{A}:=\frac{h}{u}\frac{|\nabla u|}{|\nabla h|},$
and it is easy to see $\mathcal{A}$ is positive in $\mathcal{D}$. Then (7.13) becomes $\mathcal{A}^{p}-p\mathcal{A}+(p-1)=0$, which is an equation of the variable $\mathcal{A}$ of order $p>1$. It is clear that $\mathcal{A}=1$ is a solution to the equation $\mathcal{A}^{p}-p\mathcal{A}+(p-1)=0$. However, one can see that the function $\mathcal{A}\rightarrow\mathcal{A}^{p}$ is strictly convex on $(0,\infty)$, which implies that $\mathcal{A}=1$ would be the unique solution. Hence, one has $\mathcal{A}=1$ a.e. in $\mathcal{D}$, i.e.
$h|\nabla u|=u|\nabla h|\quad\mathrm{a.e.~in}~\mathcal{D}.$ (7.14)
On $\Omega\setminus\mathcal{D}$, $|\nabla u||\nabla h|=0$. If there exists $x\in\Omega\setminus\mathcal{D}$ such that $|\nabla h(x)|=0$, then it follows from (7.10) that $|\nabla u(x)|^{p}=0$. Similarly, if there exists $x\in\Omega\setminus\mathcal{D}$ such that $|\nabla u(x)|=0$, then it follows from (7.10) that $(p-1)\frac{u^{p}(x)}{h^{p}(x)}|\nabla h(x)|^{p}=0$, which implies $|\nabla h(x)|=0$. Hence, one has
$|\nabla u|=|\nabla h|=0\quad\mathrm{a.e.~in}~\Omega\setminus\mathcal{D}.$ (7.15)
Combining (7.12), (7.14) and (7.15), we can obtain
$\chi=\frac{h}{u}.$
Since $u,h>0$ in $\Omega$ and $u,h\in C^{1,\alpha}_{\mathrm{loc}}(\Omega)$, one has
$\nabla(\log h-\log u)=\frac{\nabla h}{h}-\frac{\nabla u}{u}=\frac{\frac{h}{u}\cdot\nabla u}{h}-\frac{\nabla u}{u}=0,$
which implies that $\log h-\log u$ is $u$ is constant on each connected component of $\Omega$. Consequently, $\chi=h/u=\beta$ for some constant $\beta>0$ should be a constant function. Our proof is finished. ∎
Remark 7.2.
From the above proof of Lemma 7.1, one easily knows that its assertion is valid without requiring that the potential function $\phi$ has Property 1.
Now, we can prove:
Theorem 7.3.
Assume that an $n$-dimensional ($n\geq 2$) complete Riemannian manifold $M^{n}$ has a radial sectional curvature upper bound $\kappa(s)$ w.r.t. $q\in M^{n}$, and assume that the potential function $\phi$ has Property 1. Then for $r_{0}<\min\{\mathrm{inj}(q),l\}$ with $\mathrm{inj}(q)$ defined as in (2.5), one has
$\lambda^{\phi}_{1,p}(B(q,r_{0}))\geq\lambda^{\phi}_{1,p}\left(\mathscr{B}_{n}(q^{+},r_{0})\right),$ (7.16)
and the equality holds if and only if $B(q,r_{0})$ is isometric to $\mathscr{B}_{n}(q^{+},r_{0})$.
Proof.
Let $\psi(s):[0,r_{0}]\rightarrow[0,\infty)$ be a nonnegative radial eigenfunction of $\lambda_{1,\phi}\left(\mathscr{B}_{n}(q^{+},r_{0})\right)$ of $M^{+}$. By Lemma 6.5, one knows that the eigenfunction $\psi$ satisfies (6.8), with $\psi^{\prime}(0)=\psi(r_{0})=0$, $\psi\geq 0$ on $0,r_{0})$, and $\frac{d\psi}{ds}<0$ on $(0,r_{0})$. Define a function $\mathcal{F}:\overline{B(q,r_{0})}\rightarrow[0,\infty)$ by
$\mathcal{F}(\exp_{q}(s\xi))=\psi(s)$
for $(s,\xi)\in[0,r_{0})\times S_{q}^{n-1}$. Then by a straightforward calculation in geodesic spherical coordinates around $q$ and using the characterization (2.19), one can get
$\frac{\Delta_{p,\phi}\mathcal{F}}{\mathcal{F}^{p-1}}(\exp_{q}(s\xi))$ $\leq$ $\frac{1}{\psi^{p-1}}\left\{(p-1)|\psi^{\prime}(s)|^{p-2}\psi^{\prime\prime}(s)+\left[(n-1)\frac{f^{\prime}(s)}{f(s)}-\phi^{\prime}(s)\right]|\psi^{\prime}(s)|^{p-2}\psi^{\prime}(s)\right\}$
$=$ $-\lambda^{\phi}_{1,p}\left(\mathscr{B}_{n}(q^{+},r_{0})\right),$
which implies
$\inf\limits_{\Omega}\left\{-\frac{\Delta_{p,\phi}\mathcal{F}}{\mathcal{F}^{p-1}}\right\}\geq\lambda^{\phi}_{1,p}\left(\mathscr{B}_{n}(q^{+},r_{0})\right).$
Applying Lemma 7.1, one has
$\lambda^{\phi}_{1,p}(B(q,r_{0}))\geq\inf\limits_{B(q,r_{0})}\left\{-\frac{\Delta_{p,\phi}\mathcal{F}}{\mathcal{F}^{p-1}}\right\}\geq\lambda^{\phi}_{1,p}\left(\mathscr{B}_{n}(q^{+},r_{0})\right),$
which implies (7.16) directly. When the equality in (7.16) holds, the isometric rigidity between $B(q,r_{0})$ and $\mathscr{B}_{n}(q^{+},r_{0})$ can be obtained by directly using Theorem 2.7. ∎
Remark 7.4.
Clearly, Theorem 1.18 exactly corresponds to the case $p=2$ of Theorem 7.3.
Remark 7.5.
If $\phi=const.$, then (7.16) in Theorem 7.3 would reduce to
$\lambda_{1,p}(B(q,r_{0}))\geq\lambda_{1,p}\left(\mathscr{B}_{n}(q^{+},r_{0})\right),$ (7.17)
and the equality holds if and only if $B(q,r_{0})$ is isometric to $\mathscr{B}_{n}(q^{+},r_{0})$. If furthermore the curvature assumption was strengthened to be that the radial sectional curvature $\mathscr{K}(\cdot,\cdot)$ has an upper bound $K$ for some constant $K$, then the eigenvalue comparison (7.17) would reduce to
$\lambda_{1,p}(B(q,r_{0}))\geq\lambda_{1,p}\left(\mathcal{B}_{n}(K,r_{0})\right),$ (7.18)
and the equality holds if and only if $B(q,r_{0})$ is isometric to $\mathcal{B}_{n}(K,r_{0})$. This eigenvalue comparison (7.18) and related rigidity result have been shown in [64, Theorem 1.4] very recently.
Acknowledgments
This research was supported in part by the NSF of China (Grant Nos. 11801496 and 11926352), the Fok Ying-Tung Education Foundation (China), the Key Project of Jiangxi Provincial Natural Science Foundation (Grant No. 20252BAC250003), Hubei Key Laboratory of Applied Mathematics (Hubei University), and Key Laboratory of Intelligent Sensing System and Security (Hubei University), Ministry of Education.